### v1

# AdaSCALE: Adaptive Scaling for OOD Detection

Abstract
The ability of the deep learning model to recognize when a sample falls outside its learned distribution is critical for safe and reliable deployment. Recent state-of-the-art out-of-distribution (OOD) detection methods leverage activation shaping to improve the separation between in-distribution (ID) and OOD inputs. These approaches resort to sample-specific scaling but apply a static percentile threshold across all samples regardless of their nature, resulting in suboptimal ID-OOD separability. In this work, we propose AdaSCALE, an adaptive scaling procedure that dynamically adjusts the percentile threshold based on a sample’s estimated OOD likelihood. This estimation leverages our key observation: OOD samples exhibit significantly more pronounced activation shifts at high-magnitude activations under minor perturbation compared to ID samples. AdaSCALE enables stronger scaling for likely ID samples and weaker scaling for likely OOD samples, yielding highly separable energy scores. Our approach achieves state-of-the-art OOD detection performance, outperforming the latest rival OptFS by 14.94% in near-OOD and 21.67% in far-OOD datasets in average FPR@95 metric on the ImageNet-1k benchmark across eight diverse architectures. The code is available at: https://github.com/sudarshanregmi/AdaSCALE/
1 Introduction
The reliable deployment of deep learning models hinges on their ability to handle previously unseen inputs, a task commonly known as OOD detection. One critical application is in medical diagnosis, where a model trained on common diseases should be able to flag inputs representing unknown conditions as potential outliers, requiring further review by clinicians. OOD detection primarily involves identifying semantic shifts, with robustness to covariate shifts being a highly desirable characteristic [58, 2]. As modern deep learning models scale in both data and parameter counts, effective OOD detection within large-scale settings is critical. Given the difficulties of iterating on large models, post-hoc approaches that preserve ID accuracy are generally preferred.
A variety of post-hoc approaches have emerged, broadly categorized by where they operate. One class of methods focuses on computing OOD scores directly in the output space [16, 30, 33, 20, 7, 64], while another operates in the activation space [27, 48, 44, 40]. Finally, a more recent line of research also explores a hybrid approach [50, 24], combining information from both spaces. The efficacy of many high-performing methods relies on either accurate computation of ID statistics [47, 25, 55, 46, 26] or retention of training data statistics [48, 40]. However, as retaining full access to training data becomes increasingly impractical in large-scale settings, methods that operate effectively with minimal ID samples without performance degradation are particularly valuable for practical applications.
Alleviating the dependence on ID training data/statistics, recent state-of-the-art post-hoc approaches center around the concept of “fixed scaling.” ASH [7] prunes and scales activations on a per-sample basis. SCALE [54], the direct successor of ASH, critiques pruning and focuses purely on scaling, which improves OOD detection without accuracy degradation. LTS [8] extends this concept by directly scaling logits instead, using post-ReLU activations. These methods leverage a key insight: scaling based on the relative strength of a sample’s top-k activations (with respect to the entire activations) produces highly separable ID-OOD energy scores. However, although such approaches provide sample-specific scaling factors, the scaling mechanism remains uniform across all samples as the percentile threshold and thereby is fixed, as shown in Figure 1. This static approach is inherently limiting for optimal ID-OOD separation while also failing to leverage even minimal ID data, which could be reasonably practical in most deployment scenarios.
We hypothesize that designing an adaptive scaling procedure based on each sample’s predetermined OOD likelihood offers greater control for enhancing ID-OOD separability. Specifically, this mechanism should assign smaller scaling factors for samples with high OOD likelihood to yield lower energy scores and larger scaling factors for probable ID samples to yield higher energy scores. To achieve this, we propose a heuristic for predetermining OOD likelihood based on a key observation in activation space: minor perturbations applied to OOD samples induce significantly more pronounced shifts in their top-k activations compared to ID samples. Consequently, samples exhibiting substantial activation shifts are assigned lower scaling factors, while those with minimal shifts receive higher scaling factors. This adaptive scaling mechanism can be applied in either logit or activation space. Our method, AdaSCALE, achieves state-of-the-art performance, delivering significant improvements in OOD detection while requiring only minimal ID samples.
We conduct an extensive evaluation across 8 architectures on ImageNet-1k and 2 architectures on CIFAR benchmarks, demonstrating the substantial effectiveness of AdaSCALE. For instance, AdaSCALE surpasses the average performance of the best-generalizing method, OptFS [64], by for near-OOD detection and for far-OOD detection in terms of FPR@95 / AUROC, on the ImageNet-1k benchmark across eight architectures. Furthermore, AdaSCALE outperforms the best-performing method, SCALE [54], when evaluated on the ResNet-50 architecture, achieving performance gains of for near-OOD and for far-OOD detection. Additionally, AdaSCALE consistently demonstrates superiority in full-spectrum OOD detection [58]. Our key contributions are summarized as follows:
We reveal that OOD inputs exhibit significantly more pronounced shifts in top-k activations under minor perturbations compared to ID inputs.
Leveraging this observation, we propose a novel post-hoc OOD detection method utilizing adaptive scaling that attains state-of-the-art OOD detection.
We demonstrate state-of-the-art generalization of AdaSCALE via extensive evaluations across many setups.
2 Related Works
Post-hoc methods.
Early research on OOD detection primarily focused on designing scoring functions based on logit information [16, 30, 33, 20, 34]. While these methods leveraged logit-based scores, alternative approaches have explored gradient-based information, such as GradNorm [23], GradOrth [4], GAIA [5], and Greg-OOD [45]. Given the limited dimensionality of the logit space, which may not encapsulate sufficient information for OOD detection, subsequent studies have investigated activation-space-based methods. These approaches exploit the high-dimensional activations, leading to both parametric techniques such as MDS [27], MDS Ensemble [27], and RMDS [44], as well as non-parametric methods such as KNN-based OOD detection [48, 39]. Recent advancements have proposed hybrid methodologies that integrate parametric and non-parametric techniques to improve robustness. For instance, ComboOOD [40] combines these paradigms to enhance near-OOD detection performance. Similarly, VIM [50] employs a combination of logit-based and distance-based metrics. However, reliance of such approaches on ID statistics [46, 38, 59] can become a constraint, hindering scalability and practical deployment in real-world applications. To mitigate computational challenges for real-world deployment, recent methods, such as FDBD [32] and NCI [31], have focused on enhancing efficiency. Recent advances, such as WeiPer [14], explore class-direction perturbations, while NECO [1] examines connections to neural collapse phenomena.
Activation-shaping post-hoc methods.
A seminal work in OOD detection, ReAct [47], identified abnormally high activation patterns in OOD samples and proposed clipping extreme activations. This approach has been further generalized by BFAct [25] and VRA [55], which extend activation clipping for enhanced effectiveness. Additionally, BATS [66] refines activation distributions by aligning them with their respective typical sets, while LAPS [15] enhances this strategy by incorporating channel-aware typical sets. Inspired by activation clipping, another line of research explores activation “scaling” as a means to improve OOD detection. ASH [7] introduces a method to compute a scaling factor as a function of the activation itself, pruning and rescaling activations to enhance the separation of energy scores between ID and OOD samples. However, this approach results in a slight degradation in ID classification accuracy. In response, SCALE [54] observes that pruning adversely affects performance and thus eliminates it, leading to improved OOD detection while preserving ID accuracy. SCALE currently represents the state-of-the-art method for ResNet-50-based OOD detection. Despite their efficacy, these activation-based methods exhibit limited generalization across diverse architectures. To address this issue, LTS [8] extends SCALE by computing scaling factors using post-ReLU activations and applying them directly to logits rather than activations. Our work builds on this line of work, introducing the adaptive scaling mechanism. ATS [26] argues that relying solely on final-layer activations may result in the loss of critical information beneficial for OOD detection and proposes to leverage intermediate-layer activations too. However, its efficacy is contingent upon the availability of a large number of training samples, whereas our approach attains state-of-the-art performance while utilizing a minimal number of ID samples. A newly proposed method OptFS [64] introduces a piecewise constant shaping function with the goal of generalization across diverse architectures in large-scale settings, while our work exhibits superior generalization extending to small-scale settings too.
Training methods.
The training methods incorporate adjustments during training to enhance the ID-OOD differentiating characteristics. They either make architectural adjustments [6, 18, 22], apply enhanced data augmentations [53, 19, 21], or make simple training modifications [52, 42, 62]. More recent methods have adopted contrastive learning in the context of OOD detection [36, 43, 35, 67]. Moreover, some approaches also either utilize external real outliers [17, 60, 11, 65] or synthesize virtual outliers either in image space [10, 41, 51, 13, 63, 29, 3, 37] or in feature space [9, 49, 12, 28]. However, training strategies incur additional computational cost and potentially inferior performance to post-hoc methods in large-scale setups [57].
3 Preliminaries
Let denote the input space and denote the label space, where is the number of classes. We consider a multi-class classification setting where a classifier is trained on ID data drawn from an underlying joint distribution , where and . The ID training dataset is denoted as , where is the number of training samples and .
The classifier is composed of a feature extractor , and a classifier . The feature extractor maps an input to a feature vector , where and the classifier then maps this feature vector to a logit vector . We refer to individual dimensions of the feature vector as activations, denoted by for the -th dimension. The classifier is trained on to minimize the empirical risk:
where is a loss function, such as cross-entropy loss. During inference, the model may encounter data points drawn from a different distribution, denoted as , which is referred to as OOD data. The OOD detection problem aims to identify whether a given input is drawn from marginal distribution or from . Hence, the goal is to design a scoring function that assigns a scalar score to each input , reflecting its likelihood of being an OOD sample. A higher score typically indicates a higher probability of the input being OOD. A threshold is used to classify an input as either ID or OOD:
4 Method
In this section, we introduce AdaSCALE, a novel post-processing approach that dynamically adapts the scaling mechanism based on each sample’s estimated OOD likelihood. We first present our key empirical observations regarding activation behavior under minor perturbations, building upon insights from ReAct [47]. Next, we revisit and analyze the core principle underlying recent state-of-the-art approaches. Finally, we detail our proposed adaptive scaling mechanism that leverages these observations to achieve superior OOD detection performance.
4.1 Observations in Activation Space
A seminal work ReAct [47] demonstrated that OOD samples often induce abnormally high activations within neural networks. We extend this finding with an important observation: the positions of such high activations in OOD samples are relatively unstable under minor perturbations compared to ID samples. This instability provides a valuable signal for distinguishing OOD samples from ID samples. Below, we formalize this observation and our methodology.
4.1.1 Perturbation Mechanism
Let be an input image with input channels, height, and width. We denote channel value at position as . To identify channel values for perturbation, we employ pixel attribution that quantifies each input element’s influence on the model’s prediction. An attribution function, , assigns a score to each channel value, with lower absolute scores indicating less influence. We select of channel value indices with lowest absolute attribution scores, forming the set . We use a gradient-based attribution:
where is the predicted class index. To create a perturbed input, we select a subset containing of channel values to perturb. The perturbed image is obtained as:
where is perturbation magnitude.
4.1.2 Activation Shift as OOD Indicator
After obtaining the perturbed input , we compute its activation . We define the activation shift as the absolute element-wise difference between the original activation and the perturbed activation:
Figure 3 illustrates the key insight of our approach: activation shift at extreme (high-magnitude) activations is consistently more pronounced in OOD samples compared to ID samples. This behavior can be understood intuitively: ID samples activate network features in a stable, predictable manner reflecting learned patterns, while OOD samples trigger less stable, more arbitrary high activations that shift significantly under perturbation. Based on this observation, we propose using activation shift at the top- highest activations as a metric to estimate OOD likelihood of a sample:
where returns the indices of the highest values in .
As evidenced by ratio shown in Figure 4, the statistic generally assigns higher values to diverse OOD samples than ID samples. However, the high variance of metric (Figure 3) suggests the possibility of overoptimistic estimations. To address this issue, we introduce a correction term that exhibits an opposing behavior: it tends to be higher for ID samples than for OOD samples. Figure 5 shows that the perturbed activations of ID samples tend to be higher than those of OOD samples, especially in high-activation regions. We leverage this complementary signal by defining:
where is a hyperparameter denoting the number of considered activations. We refine our OOD quantification by combining both metrics, weighted by a hyperparameter :
If be the set of values on ID validation samples, we could transform any into a normalized probability scale by constructing empirical cumulative distribution function (eCDF) derived from . The eCDF, denoted as , can be defined as:
where is the indicator function. A higher value of indicates a higher likelihood of the sample being OOD. Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table 7).
4.2 Revisiting Static Scaling Mechanisms
Recent state-of-the-art post-processing methods for OOD detection, including SCALE [54], ASH [7], and LTS [8], operate by scaling either the activations or logits. The scaling factor is computed as:
where denotes the percentile of all elements in activation . While this approach yields sample-specific scaling factors, it imposes a critical constraint: the percentile threshold is static and identical across all test samples, regardless of the nature of samples. We argue that this static nature limits the effectiveness of the scaling procedure and prevents optimal ID-OOD separability.
Method ResNet-50 ResNet-101 RegNet-Y-16 ResNeXt-50 DenseNet-201 EfficientNetV2-L ViT-B-16 Swin-B Average near-OOD MSP 74.23 / 60.21 71.96 / 67.25 62.22 / 80.74 73.25 / 67.86 73.44 / 67.29 72.51 / 80.76 86.72 / 68.62 87.11 / 69.82 75.18 / 70.32 MLS 74.87 / 64.55 72.05 / 71.51 62.94 / 84.66 74.11 / 71.62 75.51 / 68.91 81.44 / 79.22 93.78 / 63.64 94.80 / 64.68 78.69 / 71.10 EBO 75.32 / 64.52 72.32 / 71.54 62.80 / 84.76 74.21 / 71.61 75.85 / 68.68 82.86 / 77.15 94.37 / 59.19 95.34 / 59.79 79.13 / 69.66 ReAct 72.61 / 68.81 68.07 / 75.00 70.73 / 75.37 70.96 / 74.13 69.97 / 73.65 72.36 / 71.39 86.63 / 68.35 82.64 / 73.26 74.25 / 72.50 ASH 69.47 / 71.33 65.24 / 76.61 82.51 / 67.81 70.98 / 75.25 92.83 / 52.30 94.85 / 44.78 94.45 / 53.20 96.37 / 47.58 83.34 / 61.11 SCALE 67.76 / 74.20 63.87 / 78.60 67.09 / 82.90 70.59 / 76.20 71.56 / 73.72 89.70 / 60.12 94.48 / 56.18 88.62 / 61.47 76.71 / 70.42 BFAct 72.35 / 68.88 67.96 / 75.16 78.72 / 66.09 70.96 / 74.14 71.20 / 72.61 75.53 / 62.46 82.09 / 70.66 71.81 / 75.28 73.83 / 70.66 LTS 68.01 / 73.37 63.91 / 78.27 69.82 / 80.75 70.27 / 76.20 71.29 / 74.56 87.30 / 73.63 88.83 / 67.43 86.61 / 67.22 75.76 / 73.93 OptFS 69.66 / 70.97 65.46 / 75.83 73.53 / 75.21 69.27 / 74.84 71.74 / 72.10 72.29 / 75.29 76.55 / 72.73 76.81 / 74.06 71.91 / 73.88 AdaSCALE-A 58.98 / 78.98 57.96 / 81.68 47.91 / 89.18 64.14 / 79.96 61.28 / 79.66 53.78 / 86.94 71.87 / 73.14 73.41 / 74.48 61.17 / 80.50 AdaSCALE-L 59.84 / 78.62 56.41 / 81.86 56.13 / 87.11 62.08 / 80.18 61.75 / 80.06 54.95 / 85.77 71.99 / 73.23 72.89 / 74.58 62.00 / 80.18 far-OOD MSP 53.15 / 84.06 53.87 / 83.81 40.41 / 90.08 53.07 / 84.21 53.60 / 84.43 54.74 / 87.92 56.41 / 84.62 73.39 / 82.02 54.83 / 85.14 MLS 42.57 / 88.19 43.89 / 88.30 32.92 / 93.70 44.91 / 87.97 48.43 / 87.44 68.64 / 84.80 81.89 / 81.42 95.16 / 73.37 57.30 / 85.65 EBO 42.72 / 88.09 44.30 / 88.23 32.47 / 93.82 45.12 / 87.86 48.95 / 87.15 74.48 / 81.13 86.95 / 76.34 96.08 / 63.99 58.88 / 83.33 ReAct 30.14 / 92.98 29.89 / 93.10 45.20 / 86.17 30.06 / 92.69 30.72 / 92.65 60.05 / 75.33 59.31 / 83.65 58.86 / 84.77 43.03 / 87.67 ASH 24.69 / 94.43 26.18 / 94.06 59.65 / 83.94 29.17 / 93.47 33.50 / 92.17 96.56 / 41.57 95.98 / 52.16 98.23 / 43.20 57.99 / 74.38 SCALE 21.44 / 95.39 22.54 / 95.05 32.16 / 94.16 30.62 / 93.54 33.17 / 92.70 89.63 / 62.58 88.36 / 72.32 86.59 / 66.77 50.56 / 84.06 BFAct 29.46 / 93.01 29.43 / 93.04 58.69 / 77.22 29.71 / 92.67 32.45 / 92.29 66.72 / 65.70 51.58 / 85.77 38.99 / 88.47 42.13 / 86.02 LTS 22.20 / 95.24 23.07 / 94.94 34.99 / 93.57 30.37 / 93.49 30.92 / 93.29 86.85 / 76.30 64.37 / 84.43 85.84 / 44.80 47.33 / 84.51 OptFS 25.66 / 93.87 26.97 / 93.55 47.37 / 86.73 27.54 / 93.40 34.42 / 91.04 53.62 / 83.62 46.11 / 87.35 44.27 / 87.79 38.25 / 89.67 AdaSCALE-A 17.84 / 96.14 18.51 / 95.95 21.37 / 95.84 22.08 / 95.24 28.01 / 93.23 37.61 / 91.48 47.63 / 86.83 47.81 / 87.14 30.11 / 92.73 AdaSCALE-L 17.92 / 96.12 19.15 / 95.76 20.10 / 96.19 22.16 / 95.01 28.00 / 93.18 38.81 / 90.51 47.28 / 86.97 46.24 / 87.97 29.96 / 92.71
4.3 Proposed Approach: Adaptive Scaling
Building on our observations, we propose AdaSCALE (Adaptive SCALE), a novel approach that introduces dynamic, sample-specific adjustments to the scaling procedure. The key insight of AdaSCALE is that the percentile threshold should be a function of each test sample’s estimated OOD likelihood rather than a fixed value. The scaling factor increases as the percentile threshold rises (i.e., when more activations are excluded from the denominator in Equation 10). For optimal ID-OOD separation, we must scale ID samples more strongly than OOD samples, requiring a higher percentile for ID samples. Leveraging as an OOD likelihood estimator, we can now define an adaptive percentile threshold as:
where and are hyperparameters that define the minimum and maximum limits of percentile threshold. It ensures that samples with lower OOD likelihood (i.e., likely ID samples) receive higher percentile thresholds, resulting in stronger scaling.
Algorithm 1 summarizes our adaptive scaling factor computation procedure. We implement two variants: AdaSCALE-A scales activations as , similar to ASH [7] and SCALE [54]. AdaSCALE-L scales logits as , similar to LTS [8]. For OOD detection, we use the energy score on (directly or indirectly) scaled logits, with higher values indicating higher ID likelihood. This approach enables per-sample dynamic scaling, as outlined in Figure 2.
5 Experiments
We use pre-trained models provided by PyTorch for ImageNet-1k experiments. For CIFAR experiments, we train three models per network using the standard cross-entropy loss and report the mean results across these three independent trials. The evaluation setup is provided in Table 2.
Conventional OOD detection ID datasets Near-OOD Far-OOD Network CIFAR-10/100 CIFAR-100/10 MNIST, SVHN, WRN-28-10, DenseNet-101 TIN Textures, Places365 ImageNet-1k SSB-Hard iNaturalist, Places, EfficientNetV2-L, ResNet-101 NINCO OpenImage-O DenseNet-201, ViT-B-16 ImageNet-O Textures ResNet-50, ResNeXt-50 Places RegNet-Y-16, Swin-B Covariate shifted datasets for full spectrum OOD detection ImageNet-1k ImageNet-C, ImageNet-R, ImageNet-V2, ImageNet-ES
Metrics. We use two commonly used OOD Detection metrics: Area Under Receiver-Operator Characteristics (AUROC) and False Positive Rate at 95% True Positive Rate (FPR@95), where a higher AUROC and lower FPR@95 indicates better OOD detection performance.
Baselines. We consider the following post-hoc methods: MSP [16], ODIN [30], EBO [33], ReAct [47], MLS [20], ASH [7], SCALE [54], BFAct [25], LTS [8], OptFS [64]. Currently, SCALE is the best-performing method (with ResNet-50), while OptFS is the best-generalizing method.
Hyperparameters. The hyperparameters are determined via automatic parameter search provided by OpenOOD [57, 61]. The best results are bold, and the second-best results are underlined across all results. We refer readers to Appendices A, B, C D, and E for complete details and studies.
5.1 Empirical Results
ImageNet-1k benchmark:
The ImageNet-1k benchmark is crucial for OOD evaluation, as large-scale datasets offer a comprehensive representation of real-world scenarios. Consequently, post-hoc methods are particularly valuable when training data is inaccessible and retraining is computationally prohibitive. We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table 1. AdaSCALE demonstrates consistently strong performance across all architectures compared to existing methods. Specifically, it surpasses the best-generalizing method, OptFS, by in the FPR@95/AUROC metric for near-OOD detection across all architectures. Additionally, it outperforms the best-performing method, SCALE (on ResNet-50), by in the same metric. A closer observation reveals that while OptFS excels in architectures such as EfficientNet, ViT-B-16, and Swin-B, scaling baselines perform comparably or even better in architectures like ResNet-50, ResNet-101, RegNet-Y-16, and DenseNet-201. In contrast, AdaSCALE-A achieves the best performance in near-OOD detection across all architectures, except for Swin-B, where BFAct performs optimally. Furthermore, the effectiveness of AdaSCALE extends beyond the near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% in the FPR@95 metric compared to OptFS.
Method Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC MSP 86.75 50.13 74.28 65.67 MLS 88.93 49.48 76.33 65.06 EBO 89.11 49.08 77.21 63.44 ReAct 87.22 51.38 67.23 69.53 ASH 87.01 52.02 72.36 65.70 SCALE 86.75 52.27 69.36 68.97 BFAct 87.12 51.14 66.13 69.69 LTS 86.46 53.29 66.76 71.63 OptFS 85.83 52.17 63.32 71.44 AdaSCALE-A 81.34 55.03 58.87 72.41 AdaSCALE-L 81.62 55.14 59.19 72.85
Full-Spectrum OOD Detection:
Full-spectrum OOD (FS-OOD) detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs. We present FS-OOD detection results in Table 3. We can observe that this is a highly challenging task, as covariate-shifted ID datasets cause a significant performance drop for all methods compared to the conventional case. Despite this, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in the FPR@95 metric for full-spectrum OOD detection across both near- and far-OOD datasets.
ISH training:
Apart from enhancing the prior postprocessor ASH [7], SCALE [56] introduces a training regularization to emphasize samples with more distinct ID characteristics. We assess the performance of each method in ResNet-50 and ResNet-101 model following this regularization in Table 4. The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by and in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively. Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in a full-spectrum setting. Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%. These findings also seem to generalize to the ResNet-101 network, as shown in Table 4.
Method OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD ResNet-50 MSP 74.07 / 62.16 51.13 / 84.64 87.52 / 40.36 74.41 / 61.52 MLS 74.38 / 66.43 41.57 / 88.90 88.89 / 39.49 71.53 / 61.69 EBO 74.68 / 66.46 41.85 / 88.83 89.05 / 39.18 71.77 / 61.11 ReAct 71.98 / 70.81 28.76 / 93.49 87.78 / 43.88 61.87 / 71.64 ASH 67.99 / 73.46 23.88 / 94.67 85.74 / 45.29 57.81 / 72.81 SCALE 65.68 / 76.41 20.77 / 95.62 84.31 / 48.40 54.48 / 74.79 BFAct 71.59 / 70.85 28.38 / 93.50 87.51 / 43.95 61.39 / 71.43 LTS 66.32 / 75.03 22.07 / 95.28 85.08 / 46.16 57.11 / 73.06 OptFS 67.71 / 73.03 24.65 / 94.18 85.38 / 45.91 57.09 / 72.95 AdaSCALE-A 57.43 / 80.86 16.52 / 96.46 80.85 / 51.24 51.57 / 75.63 AdaSCALE-L 56.83 / 80.81 17.62 / 96.22 80.97 / 50.59 53.43 / 74.62 ResNet-101 MSP 71.39 / 68.31 51.00 / 84.81 85.70 / 45.72 73.69 / 62.79 MLS 72.32 / 71.94 41.04 / 88.99 87.39 / 44.88 69.94 / 63.23 EBO 72.78 / 71.92 41.45 / 88.87 87.64 / 44.66 70.23 / 62.65 ReAct 67.74 / 75.74 28.53 / 93.47 85.40 / 48.83 60.50 / 72.18 ASH 66.03 / 77.79 25.21 / 94.43 83.95 / 50.57 56.91 / 73.37 SCALE 64.30 / 78.98 23.09 / 94.95 83.14 / 51.05 55.88 / 73.27 BFAct 67.53 / 75.86 28.32 / 93.40 85.11 / 48.96 60.12 / 71.83 LTS 66.32 / 75.03 22.07 / 95.28 85.08 / 46.16 57.11 / 73.06 OptFS 67.71 / 73.03 24.65 / 94.18 85.38 / 45.91 57.09 / 72.95 AdaSCALE-A 54.66 / 83.52 16.81 / 96.32 78.52 / 55.19 49.92 / 76.20 AdaSCALE-L 53.91 / 83.49 17.55 / 96.15 78.63 / 54.60 51.47 / 75.47
CIFAR benchmark:
CIFAR benchmarks are traditionally used to evaluate OOD detectors in smaller-scale settings. We compare AdaSCALE with recent state-of-the-art post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance over both architectures in Table 5. AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection. For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points. Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive. In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21. While activation-shaping methods (ReAct, ASH, SCALE, OptFS) perform remarkably well in comparison to logit-based methods in a large-scale setting (ImageNet-1k), they generally seem to underperform in small-scale settings (CIFAR). In contrast, AdaSCALE’s achieves consistently superior performance across all setups, demonstrating its robustness.
Method CIFAR-10 CIFAR-100 Near-OOD Far-OOD Near-OOD Far-OOD MSP 43.18 / 89.07 34.49 / 90.88 55.64 / 80.23 61.73 / 76.82 MLS 51.54 / 89.33 39.62 / 91.68 57.24 / 81.25 60.19 / 78.92 EBO 51.54 / 89.37 39.58 / 91.75 57.45 / 81.10 60.12 / 78.96 ReAct 49.71 / 88.59 37.32 / 92.00 63.20 / 79.58 54.78 / 80.46 ASH 78.11 / 77.97 63.12 / 83.35 80.97 / 70.09 69.38 / 79.06 SCALE 53.00 / 89.20 39.27 / 91.93 58.38 / 81.00 57.19 / 80.56 BFAct 54.90 / 88.56 43.05 / 90.66 72.26 / 74.70 57.44 / 77.63 LTS 55.71 / 88.77 41.06 / 91.74 59.98 / 80.60 80.48 / 81.79 OptFS 64.82 / 85.72 47.67 / 89.99 76.80 / 73.02 60.23 / 77.76 AdaSCALE-A 43.07 / 90.31 33.11 / 92.66 57.33 / 81.35 54.53 / 81.14 AdaSCALE-L 44.71 / 90.14 33.43 / 92.69 58.70 / 81.07 52.49 / 82.21
Accuracy:
5.2 Ablation / hyperparameter studies
Sensitivity of , , and . As discussed in Sec. 4.1.2, ID and OOD distinctions are more pronounced in high activations but diminish as more activations are included. We analyze the impact of (used in activation shift) and (used in perturbed activation) using a heatmap of validation AUROC with ResNet-50 (Figure 6(a)). The darker region indicates higher AUROC, suggesting optimal values of and for ResNet-50 model. Furthermore, the heatmap suggests that is far more critical hyperparameter than . The hyperparameter controls the weighting of in computing , the predetermined OOD likelihood. The sensitivity analysis is presented in Figure 6(b) which shows near-OOD and far-OOD detection using FPR@95. It suggests optimal FPR@95 is achieved at .
Predetermined OOD likelihood .
Adaptive scaling depends on predetermined OOD likelihood to determine the extent of scaling. We study the effect of various predetermined OOD likelihood functions on OOD detection using ResNet-50 network (ImageNet-1k) in Table 6. It clearly shows Q component of being most critical while as correction term being a relatively superior choice.
OOD Detection Near-OOD Far-OOD Q 59.43 / 78.14 19.70 / 95.73 70.39 / 74.00 21.40 / 95.31 58.97 / 78.98 17.84 / 96.14 65.11 / 76.23 19.76 / 95.67 58.91 / 78.74 18.02 / 96.08
Adaptive percentile.
Unlike SCALE which uses constant percentile, AdaSCALE uses dynamic percentile lying in range adaptive to each sample. We show the effect of various percentile limit ranges in Figure 7 in the form of a heatmap on the AUROC validation metric on ImageNet-1k benchmark. The extent of darkness in the heatmap conveys a strong performance (corresponding to highest AUROC / lowest FPR@95). The diagonal entries, representing the results of SCALE, are lighter in comparison to the rest of the cells, denoting the results of AdaSCALE. Hence, it can be observed that using adaptive percentile leads to better OOD detection in comparison to static percentile.
ID statistics.
With the rise of large models, where training data is often undisclosed or inaccessible, relying on full training ID datasets for OOD detection has become increasingly impractical. While our approach leverages ID statistics for optimal performance, we rigorously assess its effectiveness with limited data. We conduct experiments on ImageNet-1k using ID samples to compute ID statistics, where . The results in Table 7 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD 10 59.69 / 78.52 18.25 / 96.03 81.66 / 49.76 52.18 / 75.01 100 59.05 / 78.92 17.79 / 96.13 81.52 / 49.19 52.39 / 74.82 1000 58.99 / 78.95 17.86 / 96.13 81.55 / 49.48 52.24 / 74.94 5000 58.97 / 78.98 17.84 / 96.14 81.52 / 49.35 52.33 / 74.89
6 Conclusion
We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts the scaling process based on a sample’s estimated OOD likelihood. Leveraging the observation that OOD samples exhibit larger activation shifts under minor perturbations, AdaSCALE assigns stronger scaling to likely ID samples and weaker scaling to likely OOD samples, enhancing ID-OOD separability. AdaSCALE achieves state-of-the-art performance as well as generalization across architectures requiring negligibly few ID samples, making it highly practical for real-world deployment.
Supplementary Material
Appendix A Notations
Table 8 lists all the notations used in this paper.
Notation Meaning Input space. Label space. Number of classes. Number of input channels. Classifier. Underlying joint distribution of ID data. Distribution of OOD data. ID training dataset. Number of training samples. Feature extractor, parameterized by . Activation space. Classifier (mapping activations to logits), parameterized by . Logit space. Activation vector (output of ). The -th element of the activation vector . Logit vector (output of ). Loss function (e.g., cross-entropy). OOD scoring function. Threshold for classifying an input as ID or OOD. Input image. Channel value of input image at position . Height of the input image. Width of the input image. Attribution function, assigning a score to each channel value of input . Percent of channel values to perturb. Set of channel value indices with lowest absolute attribution scores. Predicted class index. Perturbation magnitude. Perturbed input image. Activation vector of the perturbed input . Activation shift vector (absolute element-wise difference between and ). , Number of highest-magnitude activations considered for and , respectively. Same as . Returns the indices that would sort the vector in descending order. Returns the maximum element of vector . , Index sets: , Sum of activation shifts for the top- activations. Correction term: sum of top- perturbed activations. Weighting factor for in the calculation. Estimated OOD likelihood. Number of ID validation samples. Set of values on the ID validation samples. Empirical cumulative distribution function (eCDF) of values. , Minimum and maximum percentile thresholds. Raw ID likelihood from eCDF Adjusted percentile threshold The -th percentile value of all elements in Scaling factor. Scaled activation vector (AdaSCALE-A). Scaled logit vector (AdaSCALE-L). Rectified Linear Unit activation function: .
Appendix B Additional studies
Image Perturbation.
A sufficiently small perturbation, as discussed in Sec. 4.3, is used to make distinctive observations in ID vs OOD. We now systematically investigate the impact of the extent and nature of perturbation on pre-determining OOD likelihood which is in-turn responsible for OOD detection. We present the results in Table 9. It is clearly evident that perturbing regions with trivial intensity value leads to better OOD detection. We can infer that with such perturbation, acts as a more reliable proxy for OOD likelihood. Another key takeaway is that randomly perturbing small regions can also lead to strong performance. This finding highlights a more efficient approach to OOD detection, relying solely on minor random perturbations rather than requiring a predetermination of trivial pixels. Furthermore, the sensitivity study of presented at Table 10 suggests the optimal value of to be around 0.5.
Region type OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD Random 61.73 / 78.15 19.44 / 95.74 83.19 / 48.59 53.45 / 74.10 59.97 / 78.67 18.14 / 96.06 81.92 / 49.19 52.35 / 74.84 60.27 / 78.02 18.45 / 96.00 82.07 / 48.62 52.84 / 74.70 62.81 / 76.27 19.95 / 95.70 83.40 / 46.55 54.94 / 73.42 Trivial 61.77 / 78.29 19.28 / 95.77 82.92 / 48.86 53.34 / 74.24 58.97 / 78.98 17.84 / 96.14 81.52 / 49.35 52.33 / 74.89 60.24 / 78.17 17.94 / 96.08 82.19 / 48.59 52.59 / 74.77 66.43 / 74.10 21.58 / 95.29 85.26 / 44.86 56.56 / 72.82 Significant 69.43 / 75.13 22.62 / 95.17 85.59 / 48.32 53.76 / 75.23 67.31 / 75.78 21.24 / 95.44 85.07 / 48.40 53.26 / 75.63 65.65 / 76.20 20.39 / 95.62 84.36 / 48.33 53.17 / 75.55 64.54 / 75.39 20.48 / 95.61 83.78 / 47.07 54.11 / 74.54 All 67.37 / 73.08 22.93 / 95.01 85.66 / 44.00 57.80 / 72.19
Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC 0.1 63.76 77.50 19.26 95.85 0.5 58.97 78.98 17.84 96.14 1.0 61.60 76.96 19.31 95.84
Latency.
AdaSCALE incurs extra computational overhead in comparison to its preceding works [7, 54, 8]. It uses an extra forward pass to obtain perturbed activation . On top of that, top-k operations (time complexity: , where D is dimension of activation space) on and are employed for pre-determining OOD likelihood. Furthermore, comparing variable percentile to fixed percentile for scaling in Table 11 across 10000 trials, we observe variable percentile inducing relatively higher latency. The latency ratio decreases as activation space with a higher dimension is considered. All experiments are performed using NVIDIA RTX 4090 GPUs.
Fixed percentile (SCALE) 33 40 45 48 54 Variable percentile (AdaSCALE) 152 149 155 152 164 Latency ratio (AdaSCALE / SCALE) 4.66 3.76 3.42 3.14 3.02
Appendix C Hyperparameters
All hyperparameters are determined with respect to the AUROC metric using automatic parameter search of OpenOOD [57, 61]. Although AdaSCALE may appear to require numerous hyperparameters, our findings indicate that setting to consistently yields near-optimal performance across all setups. Consequently, it can be inferred that only the hyperparameters and need to be appropriately tuned for any new architecture for near-optimal performance.
| Dataset | Network | Hyperparameters | ||||||
| CIFAR-10 | WideResNet-28-10 | 60 | 95 | 10 | 0.5 | |||
| DenseNet-101 | 65 | 90 | 10 | 0.5 | ||||
| CIFAR-100 | WideResNet-28-10 | 60 | 85 | 10 | 0.5 | |||
| DenseNet-101 | 70 | 80 | 10 | 0.5 | ||||
| ImageNet-1k | ResNet-50 | 80 | 85 | 10 | 0.5 | |||
| ResNet-101 | 80 | 85 | 10 | 0.5 | ||||
| RegNet-Y-16 | 60 | 90 | 10 | 0.5 | ||||
| ResNeXt-50 | 80 | 85 | 10 | 0.5 | ||||
| DenseNet-201 | 90 | 95 | 10 | 0.5 | ||||
| EfficientNetV2-L | 60 | 99 | 10 | 0.5 | ||||
| Vit-B-16 | 60 | 85 | 10 | 0.5 | ||||
| Swin-B | 90 | 99 | 10 | 0.5 | ||||
| Dataset | Network | Hyperparameters | ||||||
| CIFAR-10 | WideResNet-28-10 | 60 | 85 | 10 | 0.5 | |||
| DenseNet-101 | 70 | 85 | 10 | 0.5 | ||||
| CIFAR-100 | WideResNet-28-10 | 60 | 80 | 10 | 0.5 | |||
| DenseNet-101 | 65 | 75 | 10 | 0.5 | ||||
| ImageNet-1k | ResNet-50 | 80 | 85 | 10 | 0.5 | |||
| ResNet-101 | 70 | 80 | 10 | 0.5 | ||||
| RegNet-Y-16 | 60 | 85 | 10 | 0.5 | ||||
| ResNeXt-50 | 70 | 80 | 10 | 0.5 | ||||
| DenseNet-201 | 90 | 95 | 10 | 0.5 | ||||
| EfficientNetV2-L | 60 | 99 | 10 | 0.5 | ||||
| Vit-B-16 | 75 | 85 | 10 | 0.5 | ||||
| Swin-B | 90 | 99 | 10 | 0.5 | ||||
Appendix D CIFAR-results
D.1 WRN-28-10
CIFAR-10 benchmark Method MNIST SVHN Textures Places365 Average MSP 17.02 / 94.61 21.71 / 92.96 60.50 / 88.06 42.27 / 90.04 35.38 / 91.42 MLS 13.01 / 96.76 30.35 / 93.06 76.12 / 86.65 52.56 / 90.46 43.01 / 91.73 EBO 12.93 / 96.93 30.35 / 93.12 76.15 / 86.68 52.57 / 90.56 43.00 / 91.82 ReAct 15.50 / 96.30 34.01 / 92.47 57.76 / 88.77 57.33 / 89.66 41.15 / 91.80 ASH 50.11 / 88.80 89.90 / 74.76 95.07 / 72.91 92.22 / 70.06 81.82 / 76.63 SCALE 13.24 / 96.70 32.21 / 92.88 75.76 / 86.77 55.81 / 90.09 44.26 / 91.61 BFAct 25.79 / 94.64 43.08 / 91.10 57.16 / 88.80 61.00 / 88.32 46.75 / 90.71 LTS 14.04 / 96.60 39.85 / 92.21 76.85 / 86.43 63.13 / 89.19 48.47 / 91.11 OptFS 25.68 / 94.83 51.58 / 89.86 62.14 / 88.07 80.19 / 84.05 54.90 / 89.20 AdaSCALE-A 14.93 / 96.02 17.84 / 95.14 64.96 / 88.31 34.57 / 92.31 33.08 / 92.95 AdaSCALE-L 15.58 / 95.98 18.41 / 95.10 62.87 / 88.67 37.59 / 91.97 33.61 / 92.93 CIFAR-100 benchmark MNIST SVHN Textures Places365 Average MSP 49.79 / 78.72 56.76 / 80.70 64.49 / 76.86 56.66 / 79.96 56.92 / 79.06 MLS 46.57 / 81.43 53.08 / 83.37 64.59 / 77.65 59.70 / 79.82 55.99 / 80.57 EBO 46.41 / 81.99 52.92 / 83.77 64.58 / 77.61 59.76 / 79.60 55.92 / 80.74 ReAct 49.92 / 81.07 40.66 / 86.49 52.42 / 80.81 60.35 / 79.72 50.84 / 82.03 ASH 44.06 / 85.55 41.48 / 87.50 61.78 / 81.65 80.45 / 71.83 56.94 / 81.63 SCALE 40.65 / 84.68 48.56 / 85.56 58.45 / 80.81 60.51 / 79.90 52.04 / 82.74 BFAct 61.59 / 77.47 34.74 / 88.50 47.30 / 83.38 64.49 / 78.47 52.03 / 81.96 LTS 36.27 / 87.38 45.41 / 87.23 53.90 / 83.18 62.62 / 79.64 49.55 / 84.36 OptFS 57.61 / 79.47 37.04 / 86.43 53.02 / 80.43 70.44 / 76.77 54.53 / 80.78 AdaSCALE-A 45.18 / 81.69 36.79 / 89.20 55.93 / 81.93 56.48 / 81.55 48.59 / 83.59 AdaSCALE-L 42.13 / 83.58 32.44 / 91.02 50.87 / 84.14 57.83 / 81.51 45.82 / 85.06
Method CIFAR-10 benchmark CIFAR-100 benchmark Average CIFAR-100 TIN CIFAR-10 TIN MSP 54.13 / 88.28 42.94 / 89.93 56.83 / 80.42 48.82 / 83.39 50.68 / 85.51 MLS 67.10 / 87.72 55.91 / 89.90 58.99 / 80.98 49.27 / 84.01 57.82 / 85.65 EBO 67.04 / 87.77 55.88 / 89.97 58.97 / 80.93 49.39 / 83.95 57.82 / 85.66 ReAct 65.96 / 87.76 51.44 / 90.33 69.17 / 79.10 51.56 / 83.77 59.53 / 85.24 ASH 91.33 / 70.72 90.77 / 73.28 85.25 / 69.96 78.31 / 74.55 86.42 / 72.13 SCALE 69.52 / 87.35 59.48 / 89.53 61.30 / 80.35 51.25 / 83.65 60.39 / 85.22 BFAct 66.31 / 86.94 57.12 / 89.69 78.90 / 74.98 59.04 / 82.25 65.34 / 83.47 LTS 74.28 / 86.38 66.01 / 88.56 64.17 / 79.57 54.24 / 83.09 64.68 / 84.40 OptFS 76.36 / 84.05 66.73 / 86.56 85.40 / 75.55 64.11 / 80.99 73.15 / 79.83 AdaSCALE-A 50.60 / 89.40 42.80 / 91.13 62.21 / 79.99 47.11 / 84.98 50.68 / 86.38 AdaSCALE-L 53.98 / 89.01 45.95 / 90.77 65.41 / 79.27 48.74 / 84.75 53.52 / 85.95
D.2 DenseNet-101
CIFAR-10 benchmark Method MNIST SVHN Textures Places365 Average MSP 17.91 / 94.22 32.04 / 90.38 46.80 / 87.53 37.59 / 89.24 33.59 / 90.34 MLS 10.02 / 97.58 31.25 / 92.59 64.43 / 85.58 39.19 / 90.74 36.22 / 91.62 EBO 9.74 / 97.76 31.23 / 92.69 64.46 / 85.48 39.17 / 90.81 36.15 / 91.68 ReAct 12.60 / 97.24 34.79 / 92.02 50.41 / 88.21 36.12 / 91.35 33.48 / 92.20 ASH 9.40 / 98.12 39.42 / 91.25 70.95 / 85.39 57.90 / 85.48 44.42 / 90.06 SCALE 9.04 / 97.88 26.99 / 93.54 61.52 / 86.77 39.57 / 90.76 34.28 / 92.24 BFAct 23.59 / 94.96 42.49 / 89.00 53.74 / 87.38 37.53 / 91.09 39.34 / 90.61 LTS 8.92 / 97.97 27.16 / 93.59 59.07 / 87.09 39.47 / 90.81 33.65 / 92.37 OptFS 9.74 / 97.88 41.20 / 90.71 51.35 / 88.48 59.47 / 86.03 40.44 / 90.77 AdaSCALE-A 12.42 / 96.85 25.04 / 94.05 58.28 / 87.35 36.77 / 91.20 33.13 / 92.36 AdaSCALE-L 10.92 / 97.44 26.43 / 93.87 58.59 / 87.19 37.03 / 91.25 33.24 / 92.44 CIFAR-100 benchmark MSP 65.65 / 72.43 63.81 / 76.52 75.34 / 72.19 61.36 / 77.16 66.54 / 74.57 MLS 58.69 / 78.55 57.12 / 79.43 79.05 / 72.68 62.72 / 78.38 64.39 / 77.26 EBO 58.58 / 78.98 56.76 / 79.19 79.09 / 72.44 62.86 / 78.08 64.32 / 77.17 ReAct 62.71 / 76.37 48.48 / 81.64 64.65 / 78.62 59.00 / 78.89 58.71 / 78.88 ASH 40.69 / 88.57 48.03 / 86.24 65.24 / 83.29 73.29 / 71.88 56.81 / 82.49 SCALE 56.92 / 79.53 53.81 / 80.96 76.10 / 74.47 62.49 / 78.51 62.33 / 78.37 BFAct 73.83 / 67.19 60.01 / 75.85 69.29 / 76.41 68.15 / 73.73 67.82 / 73.29 LTS 55.33 / 80.58 51.33 / 82.04 73.06 / 75.89 62.58 / 78.36 60.58 / 79.22 OptFS 64.24 / 75.24 59.81 / 76.46 66.15 / 77.50 73.47 / 69.76 65.92 / 74.74 AdaSCALE-A 62.51 / 74.96 46.29 / 84.31 71.40 / 76.59 61.70 / 78.86 60.47 / 78.68 AdaSCALE-L 61.33 / 75.73 43.97 / 85.30 69.31 / 77.71 61.97 / 78.69 59.15 / 79.36
Method CIFAR-10 benchmark CIFAR-100 benchmark Average CIFAR-100 TIN CIFAR-10 TIN MSP 40.13 / 88.45 35.50 / 89.61 59.94 / 77.53 56.96 / 79.57 48.13 / 83.79 MLS 45.14 / 88.85 38.01 / 90.85 63.61 / 78.26 57.09 / 81.75 50.96 / 84.93 EBO 45.19 / 88.85 38.05 / 90.90 63.90 / 77.94 57.53 / 81.58 51.17 / 84.82 ReAct 44.34 / 89.19 37.10 / 91.08 70.77 / 75.06 61.30 / 80.40 53.38 / 83.93 ASH 67.78 / 82.68 62.54 / 85.18 81.65 / 65.84 78.66 / 70.01 72.66 / 75.93 SCALE 45.25 / 88.92 37.76 / 91.01 64.20 / 78.13 56.96 / 81.88 51.04 / 84.99 BFAct 52.06 / 87.70 44.09 / 89.89 79.31 / 67.39 71.79 / 74.18 61.81 / 79.79 LTS 44.89 / 89.03 37.67 / 91.10 64.66 / 77.87 56.83 / 81.87 51.01 / 84.97 OptFS 60.63 / 85.29 55.55 / 86.96 82.97 / 64.99 74.73 / 70.55 68.47 / 76.95 AdaSCALE-A 43.29 / 89.37 35.57 / 91.32 65.51 / 78.00 54.49 / 82.44 49.72 / 85.28 AdaSCALE-L 43.19 / 89.40 35.70 / 91.37 66.09 / 77.79 54.57 / 82.47 49.89 / 85.26
Appendix E ImageNet-1k results
E.1 near-OOD detection
Method SSB-Hard NINCO ImageNet-O Average MSP 74.49 / 72.09 56.88 / 79.95 91.32 / 28.60 74.23 / 60.21 MLS 76.20 / 72.51 59.44 / 80.41 88.97 / 40.73 74.87 / 64.55 EBO 76.54 / 72.08 60.58 / 79.70 88.84 / 41.78 75.32 / 64.52 REACT 77.55 / 73.03 55.82 / 81.73 84.45 / 51.67 72.61 / 68.81 ASH 73.66 / 72.89 53.05 / 83.45 81.70 / 57.67 69.47 / 71.33 SCALE 67.72 / 77.35 51.80 / 85.37 83.77 / 59.89 67.76 / 74.20 BFAct 77.20 / 73.15 55.27 / 81.88 84.57 / 51.62 72.35 / 68.88 LTS 68.46 / 77.10 51.24 / 85.33 84.33 / 57.69 68.01 / 73.37 OptFS 78.32 / 71.01 52.09 / 82.51 78.56 / 59.40 69.66 / 70.97 AdaSCALE-A 57.96 / 81.68 44.92 / 87.15 74.06 / 68.12 58.98 / 78.98 AdaSCALE-L 58.68 / 81.42 45.01 / 87.11 75.83 / 67.33 59.84 / 78.62
Method SSB-Hard NINCO ImageNet-O Average MSP 73.20 / 72.57 55.27 / 80.61 87.42 / 48.57 71.96 / 67.25 MLS 74.68 / 74.37 55.65 / 82.29 85.81 / 57.89 72.05 / 71.51 EBO 74.96 / 74.12 56.33 / 81.79 85.66 / 58.72 72.32 / 71.54 REACT 75.96 / 74.43 52.58 / 83.27 75.67 / 67.31 68.07 / 75.00 ASH 72.48 / 74.23 49.41 / 84.62 73.84 / 70.98 65.24 / 76.61 SCALE 68.47 / 77.10 49.03 / 86.20 74.09 / 72.50 63.87 / 78.60 BFAct 75.48 / 74.74 52.23 / 83.37 76.16 / 67.37 67.96 / 75.16 OptFS 76.55 / 72.29 50.89 / 83.35 68.94 / 71.85 65.46 / 75.83 AdaSCALE-A 61.00 / 80.29 46.70 / 86.99 62.05 / 78.27 56.59 / 81.85 AdaSCALE-L 61.05 / 80.41 47.77 / 86.84 60.40 / 78.35 56.41 / 81.86
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 65.35 / 78.28 | 48.48 / 86.85 | 72.82 / 77.09 | 62.22 / 80.74 |
| MLS | 62.48 / 84.83 | 42.76 / 91.56 | 83.60 / 77.58 | 62.94 / 84.66 |
| EBO | 62.10 / 85.28 | 42.49 / 91.67 | 83.82 / 77.33 | 62.80 / 84.76 |
| REACT | 73.02 / 73.17 | 59.81 / 80.91 | 79.37 / 72.02 | 70.73 / 75.37 |
| ASH | 80.58 / 67.70 | 77.23 / 71.42 | 89.71 / 64.30 | 82.51 / 67.81 |
| SCALE | 66.98 / 82.35 | 49.84 / 89.93 | 84.44 / 76.43 | 67.09 / 82.90 |
| BFAct | 79.40 / 64.39 | 73.98 / 70.35 | 82.76 / 63.54 | 78.72 / 66.09 |
| LTS | 69.52 / 79.78 | 55.38 / 87.71 | 84.55 / 74.78 | 69.82 / 80.75 |
| OptFS | 79.59 / 69.47 | 63.97 / 80.36 | 77.03 / 75.79 | 73.53 / 75.21 |
| AdaSCALE-A | 54.50 / 87.21 | 31.50 / 93.50 | 57.75 / 86.83 | 47.91 / 89.18 |
| AdaSCALE-L | 62.61 / 84.60 | 47.84 / 90.13 | 57.94 / 86.61 | 56.13 / 87.11 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 73.04 / 73.28 | 57.90 / 80.86 | 88.81 / 49.43 | 73.25 / 67.86 |
| MLS | 74.68 / 75.06 | 60.79 / 81.91 | 86.87 / 57.87 | 74.11 / 71.61 |
| EBO | 74.90 / 74.89 | 60.96 / 81.44 | 86.76 / 58.49 | 74.21 / 71.61 |
| REACT | 75.54 / 74.51 | 57.29 / 82.50 | 80.03 / 65.37 | 70.95 / 74.13 |
| ASH | 70.72 / 76.64 | 58.40 / 83.49 | 83.84 / 65.63 | 70.99 / 75.25 |
| SCALE | 67.77 / 79.73 | 56.87 / 85.39 | 87.15 / 63.48 | 70.60 / 76.20 |
| BFAct | 75.36 / 74.65 | 57.65 / 82.46 | 79.86 / 65.30 | 70.96 / 74.14 |
| LTS | 68.26 / 79.36 | 56.35 / 85.39 | 86.22 / 63.85 | 70.28 / 76.20 |
| OptFS | 75.62 / 73.82 | 57.07 / 82.37 | 75.13 / 68.33 | 69.27 / 74.84 |
| AdaSCALE-A | 61.03 / 81.86 | 50.80 / 86.54 | 80.57 / 71.48 | 64.13 / 79.96 |
| AdaSCALE-L | 61.57 / 81.11 | 48.78 / 86.40 | 75.88 / 73.02 | 62.08 / 80.18 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 74.43 / 72.23 | 56.69 / 80.85 | 89.18 / 48.80 | 73.44 / 67.29 |
| MLS | 76.62 / 72.48 | 60.14 / 80.91 | 89.78 / 53.34 | 75.51 / 68.91 |
| EBO | 76.92 / 72.00 | 60.88 / 80.01 | 89.75 / 54.03 | 75.85 / 68.68 |
| ReAct | 78.62 / 70.93 | 57.51 / 81.19 | 73.78 / 68.83 | 69.97 / 73.65 |
| ASH | 78.80 / 68.71 | 63.84 / 79.45 | 80.07 / 68.19 | 74.24 / 72.12 |
| SCALE | 73.64 / 74.43 | 56.90 / 83.80 | 84.14 / 62.92 | 71.56 / 73.72 |
| BFAct | 81.57 / 67.52 | 65.10 / 77.38 | 66.93 / 72.93 | 71.20 / 72.61 |
| LTS | 73.46 / 74.36 | 57.54 / 83.79 | 82.87 / 65.52 | 71.29 / 74.56 |
| OptFS | 82.76 / 65.38 | 63.26 / 78.12 | 69.21 / 72.79 | 71.74 / 72.10 |
| AdaSCALE-A | 68.46 / 77.10 | 56.66 / 84.32 | 58.72 / 77.55 | 61.28 / 79.66 |
| AdaSCALE-L | 68.97 / 76.85 | 57.96 / 83.92 | 58.30 / 79.41 | 61.75 / 80.06 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 81.28 / 75.03 | 57.97 / 86.70 | 78.26 / 80.53 | 72.51 / 80.76 |
| MLS | 84.74 / 73.50 | 72.88 / 84.83 | 86.71 / 79.32 | 81.44 / 79.22 |
| EBO | 85.27 / 71.58 | 75.81 / 82.07 | 87.49 / 77.81 | 82.86 / 77.15 |
| ReAct | 74.29 / 70.63 | 71.93 / 70.92 | 70.86 / 72.63 | 72.36 / 71.39 |
| ASH | 94.82 / 46.73 | 96.44 / 37.79 | 93.30 / 49.81 | 94.85 / 44.78 |
| SCALE | 90.16 / 57.07 | 89.93 / 59.69 | 89.03 / 63.60 | 89.70 / 60.12 |
| BFAct | 75.36 / 63.66 | 77.03 / 59.56 | 74.19 / 64.18 | 75.53 / 62.46 |
| LTS | 88.43 / 68.29 | 86.68 / 75.87 | 86.78 / 76.73 | 87.30 / 73.63 |
| OptFS | 74.68 / 73.83 | 70.24 / 76.18 | 71.94 / 75.86 | 72.29 / 75.29 |
| AdaSCALE-A | 60.84 / 83.48 | 47.45 / 89.47 | 53.04 / 87.87 | 53.78 / 86.94 |
| AdaSCALE-L | 53.56 / 85.00 | 58.55 / 84.75 | 52.72 / 87.58 | 54.95 / 85.77 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 86.41 / 68.94 | 77.28 / 78.11 | 96.48 / 58.81 | 86.72 / 68.62 |
| MLS | 91.52 / 64.20 | 92.98 / 72.40 | 96.84 / 54.33 | 93.78 / 63.64 |
| EBO | 92.24 / 58.80 | 94.14 / 66.02 | 96.74 / 52.74 | 94.37 / 59.19 |
| ReAct | 90.46 / 63.10 | 78.50 / 75.43 | 90.94 / 66.53 | 86.63 / 68.35 |
| ASH | 93.50 / 53.90 | 95.37 / 52.51 | 94.47 / 53.19 | 94.45 / 53.20 |
| SCALE | 92.37 / 56.55 | 94.62 / 61.52 | 96.44 / 50.47 | 94.48 / 56.18 |
| BFAct | 89.81 / 64.16 | 71.37 / 78.06 | 85.09 / 69.75 | 82.09 / 70.66 |
| LTS | 91.42 / 64.35 | 82.63 / 75.48 | 92.42 / 62.46 | 88.83 / 67.43 |
| OptFS | 87.98 / 66.30 | 64.24 / 80.46 | 77.43 / 71.43 | 76.55 / 72.73 |
| AdaSCALE-A | 85.89 / 66.57 | 61.92 / 80.47 | 67.81 / 72.37 | 71.87 / 73.14 |
| AdaSCALE-L | 86.19 / 66.25 | 61.79 / 80.42 | 67.99 / 73.01 | 71.99 / 73.23 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 86.47 / 71.30 | 77.95 / 78.50 | 96.90 / 59.65 | 87.11 / 69.82 |
| MLS | 94.05 / 65.04 | 93.38 / 71.75 | 96.97 / 57.26 | 94.80 / 64.68 |
| EBO | 94.66 / 58.96 | 94.59 / 64.02 | 96.75 / 56.40 | 95.34 / 59.79 |
| ReAct | 89.19 / 68.70 | 68.54 / 80.16 | 90.20 / 70.93 | 82.64 / 73.26 |
| ASH | 97.15 / 45.47 | 96.64 / 47.36 | 95.32 / 49.92 | 96.37 / 47.58 |
| SCALE | 90.84 / 56.53 | 87.86 / 62.49 | 87.16 / 65.38 | 88.62 / 61.47 |
| BFAct | 84.86 / 69.41 | 61.30 / 81.10 | 69.27 / 75.34 | 71.81 / 75.28 |
| LTS | 90.36 / 64.51 | 81.02 / 74.23 | 88.44 / 62.92 | 86.61 / 67.22 |
| OptFS | 88.68 / 68.43 | 66.36 / 80.27 | 75.38 / 73.49 | 76.81 / 74.06 |
| AdaSCALE-A | 80.10 / 70.46 | 64.67 / 81.10 | 75.46 / 71.87 | 73.41 / 74.48 |
| AdaSCALE-L | 80.12 / 70.06 | 63.68 / 81.35 | 74.87 / 72.34 | 72.89 / 74.58 |
E.2 far-OOD detection
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 43.34 / 88.41 | 60.87 / 82.43 | 50.13 / 84.86 | 58.26 / 80.55 | 53.15 / 84.06 |
| MLS | 30.61 / 91.17 | 46.17 / 88.39 | 37.88 / 89.17 | 55.62 / 84.05 | 42.57 / 88.19 |
| EBO | 31.30 / 90.63 | 45.77 / 88.70 | 38.09 / 89.06 | 55.73 / 83.97 | 42.72 / 88.09 |
| ReAct | 16.72 / 96.34 | 29.64 / 92.79 | 32.58 / 91.87 | 41.62 / 90.93 | 30.14 / 92.98 |
| ASH | 14.09 / 97.06 | 15.30 / 96.90 | 29.19 / 93.26 | 40.16 / 90.48 | 24.69 / 94.43 |
| SCALE | 9.50 / 98.02 | 11.90 / 97.63 | 28.18 / 93.95 | 36.18 / 91.96 | 21.44 / 95.39 |
| BFAct | 15.94 / 96.47 | 28.43 / 92.87 | 32.66 / 91.90 | 40.83 / 90.79 | 29.46 / 93.01 |
| LTS | 10.24 / 97.87 | 13.06 / 97.42 | 27.81 / 94.01 | 37.68 / 91.65 | 22.20 / 95.24 |
| OptFS | 15.88 / 96.65 | 16.60 / 96.10 | 29.94 / 92.53 | 40.24 / 90.20 | 25.66 / 93.87 |
| AdaSCALE-A | 7.61 / 98.31 | 10.57 / 97.88 | 20.67 / 95.62 | 32.60 / 92.74 | 17.86 / 96.14 |
| AdaSCALE-L | 7.78 / 98.29 | 10.33 / 97.92 | 20.61 / 95.62 | 32.97 / 92.63 | 17.92 / 96.12 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 48.30 / 86.27 | 59.00 / 83.60 | 49.36 / 84.82 | 58.84 / 80.56 | 53.87 / 83.81 |
| MLS | 41.11 / 88.83 | 43.59 / 89.85 | 38.13 / 89.25 | 52.74 / 85.28 | 43.89 / 88.30 |
| EBO | 41.65 / 88.30 | 43.66 / 90.14 | 38.48 / 89.12 | 53.42 / 85.37 | 44.30 / 88.23 |
| ReAct | 19.86 / 95.66 | 26.94 / 93.78 | 30.18 / 92.54 | 42.58 / 90.41 | 29.89 / 93.10 |
| ASH | 19.90 / 95.68 | 13.94 / 97.32 | 27.76 / 93.63 | 43.11 / 89.59 | 26.18 / 94.06 |
| SCALE | 13.90 / 97.05 | 9.34 / 98.04 | 25.91 / 94.47 | 40.99 / 90.64 | 22.54 / 95.05 |
| BFAct | 19.60 / 95.69 | 25.79 / 93.79 | 30.18 / 92.55 | 42.14 / 90.13 | 29.43 / 93.04 |
| LTS | 15.07 / 96.83 | 10.33 / 97.89 | 25.51 / 94.52 | 41.40 / 90.53 | 23.07 / 94.94 |
| OptFS | 19.11 / 95.70 | 16.53 / 96.35 | 28.76 / 92.94 | 43.47 / 89.22 | 26.97 / 93.55 |
| AdaSCALE-A | 10.74 / 97.64 | 8.90 / 98.21 | 18.75 / 96.03 | 35.66 / 91.92 | 18.51 / 95.95 |
| AdaSCALE-L | 11.71 / 97.36 | 10.44 / 97.93 | 17.87 / 96.18 | 36.57 / 91.55 | 19.15 / 95.76 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 28.13 / 94.67 | 44.73 / 88.48 | 36.27 / 91.96 | 52.51 / 85.21 | 40.41 / 90.08 |
| MLS | 9.10 / 98.05 | 39.74 / 92.82 | 25.71 / 95.70 | 57.14 / 88.22 | 32.92 / 93.70 |
| EBO | 7.72 / 98.29 | 38.18 / 93.02 | 25.94 / 95.83 | 58.04 / 88.13 | 32.47 / 93.82 |
| ReAct | 21.24 / 94.14 | 41.20 / 87.25 | 43.46 / 89.20 | 74.92 / 74.10 | 45.20 / 86.17 |
| ASH | 48.89 / 87.39 | 45.75 / 88.79 | 70.98 / 82.52 | 72.99 / 77.06 | 59.65 / 83.94 |
| SCALE | 11.13 / 97.88 | 28.29 / 95.31 | 33.59 / 94.87 | 55.62 / 88.59 | 32.16 / 94.16 |
| BFAct | 37.88 / 86.24 | 54.87 / 77.64 | 62.53 / 79.59 | 79.46 / 65.39 | 58.69 / 77.22 |
| LTS | 14.29 / 97.52 | 25.21 / 95.72 | 43.38 / 93.53 | 57.08 / 87.51 | 34.99 / 93.57 |
| OptFS | 28.95 / 93.68 | 39.99 / 90.13 | 44.96 / 89.85 | 75.59 / 73.24 | 47.37 / 86.73 |
| AdaSCALE-A | 4.34 / 99.09 | 26.06 / 95.21 | 13.09 / 97.57 | 41.98 / 91.48 | 21.37 / 95.84 |
| AdaSCALE-L | 4.41 / 99.02 | 13.50 / 97.61 | 18.56 / 96.92 | 43.93 / 91.22 | 20.10 / 96.19 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 43.56 / 88.04 | 62.23 / 82.13 | 48.06 / 85.65 | 58.42 / 81.02 | 53.07 / 84.21 |
| MLS | 32.96 / 90.93 | 51.58 / 87.39 | 37.33 / 89.80 | 57.76 / 83.77 | 44.91 / 87.97 |
| EBO | 33.42 / 90.54 | 51.73 / 87.56 | 37.79 / 89.72 | 57.56 / 83.62 | 45.12 / 87.86 |
| ReAct | 17.64 / 95.95 | 32.86 / 91.67 | 29.82 / 92.37 | 39.92 / 90.76 | 30.06 / 92.69 |
| ASH | 17.90 / 96.22 | 23.74 / 95.18 | 30.83 / 93.13 | 44.21 / 89.35 | 29.17 / 93.47 |
| SCALE | 15.66 / 96.75 | 27.75 / 94.94 | 31.43 / 93.41 | 47.62 / 89.08 | 30.62 / 93.54 |
| BFAct | 17.40 / 95.91 | 32.00 / 91.83 | 29.53 / 92.38 | 39.89 / 90.57 | 29.71 / 92.67 |
| LTS | 16.29 / 96.63 | 26.64 / 95.07 | 30.50 / 93.50 | 48.04 / 88.78 | 30.37 / 93.49 |
| OptFS | 17.20 / 96.12 | 23.11 / 94.69 | 29.59 / 92.75 | 40.24 / 90.05 | 27.54 / 93.40 |
| AdaSCALE-A | 10.02 / 97.80 | 17.99 / 96.38 | 22.93 / 95.17 | 37.38 / 91.62 | 22.08 / 95.24 |
| AdaSCALE-L | 11.28 / 97.45 | 18.46 / 96.20 | 21.23 / 95.35 | 37.68 / 91.03 | 22.16 / 95.01 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 42.02 / 89.84 | 62.33 / 81.56 | 50.31 / 85.19 | 59.74 / 81.14 | 53.60 / 84.43 |
| MLS | 31.99 / 92.11 | 57.75 / 85.56 | 42.70 / 88.28 | 61.30 / 83.82 | 48.43 / 87.44 |
| EBO | 33.12 / 91.46 | 57.47 / 85.55 | 43.75 / 87.91 | 61.46 / 83.67 | 48.95 / 87.15 |
| ReAct | 19.41 / 95.64 | 23.86 / 94.63 | 32.54 / 91.83 | 47.06 / 88.52 | 30.72 / 92.65 |
| ASH | 21.57 / 95.47 | 21.42 / 95.56 | 41.23 / 90.19 | 49.80 / 87.45 | 33.50 / 92.17 |
| SCALE | 18.13 / 96.29 | 27.22 / 94.52 | 34.52 / 92.15 | 52.82 / 87.83 | 33.17 / 92.70 |
| BFAct | 20.64 / 95.42 | 21.70 / 95.17 | 39.76 / 89.97 | 47.72 / 88.61 | 32.45 / 92.29 |
| LTS | 15.68 / 96.71 | 22.49 / 95.81 | 34.27 / 92.37 | 51.23 / 88.26 | 30.92 / 93.29 |
| OptFS | 25.81 / 93.92 | 21.75 / 95.01 | 38.45 / 89.67 | 51.66 / 85.54 | 34.42 / 91.04 |
| AdaSCALE-A | 17.30 / 96.03 | 19.42 / 96.23 | 23.12 / 94.68 | 52.20 / 85.98 | 28.01 / 93.23 |
| AdaSCALE-L | 17.97 / 95.87 | 16.87 / 96.69 | 23.64 / 94.69 | 53.50 / 85.46 | 28.00 / 93.18 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 25.14 / 95.12 | 74.42 / 84.20 | 40.64 / 91.74 | 78.74 / 80.61 | 54.74 / 87.92 |
| MLS | 35.28 / 94.13 | 86.65 / 80.26 | 62.11 / 90.26 | 90.53 / 74.56 | 68.64 / 84.80 |
| EBO | 49.84 / 91.21 | 87.72 / 75.77 | 68.77 / 87.66 | 91.60 / 69.89 | 74.48 / 81.13 |
| ReAct | 46.44 / 80.96 | 54.56 / 77.17 | 60.79 / 78.20 | 78.39 / 64.99 | 60.05 / 75.33 |
| ASH | 96.26 / 37.76 | 95.40 / 50.98 | 97.52 / 43.19 | 97.07 / 34.34 | 96.56 / 41.57 |
| SCALE | 87.08 / 67.69 | 86.22 / 67.44 | 91.05 / 67.21 | 94.18 / 47.99 | 89.63 / 62.58 |
| BFAct | 57.31 / 69.11 | 63.43 / 67.70 | 69.30 / 67.49 | 76.86 / 58.52 | 66.72 / 65.70 |
| LTS | 79.05 / 84.72 | 86.89 / 75.39 | 88.00 / 81.53 | 93.45 / 63.56 | 86.85 / 76.30 |
| OptFS | 38.62 / 89.80 | 45.77 / 86.94 | 53.77 / 85.49 | 76.31 / 72.23 | 53.62 / 83.62 |
| AdaSCALE-A | 18.51 / 96.67 | 42.07 / 90.56 | 31.00 / 94.44 | 58.87 / 84.26 | 37.61 / 91.48 |
| AdaSCALE-L | 26.58 / 95.02 | 32.81 / 92.38 | 39.19 / 92.31 | 56.66 / 82.33 | 38.81 / 90.51 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 42.40 / 88.19 | 56.46 / 85.06 | 56.19 / 84.86 | 70.59 / 80.38 | 56.41 / 84.62 |
| MLS | 72.98 / 85.29 | 78.93 / 83.74 | 85.78 / 81.60 | 89.88 / 75.05 | 81.89 / 81.42 |
| EBO | 83.56 / 79.30 | 83.66 / 81.17 | 88.82 / 76.48 | 91.77 / 68.42 | 86.95 / 76.34 |
| ReAct | 48.22 / 86.11 | 55.87 / 86.66 | 57.68 / 84.29 | 75.48 / 77.52 | 59.31 / 83.65 |
| ASH | 97.02 / 50.62 | 98.50 / 48.53 | 94.79 / 55.51 | 93.60 / 53.97 | 95.98 / 52.16 |
| SCALE | 86.60 / 73.94 | 84.70 / 79.00 | 89.48 / 72.72 | 92.67 / 63.60 | 88.36 / 72.32 |
| BFAct | 40.56 / 87.96 | 48.65 / 88.31 | 48.24 / 86.59 | 68.86 / 80.21 | 51.58 / 85.77 |
| LTS | 50.42 / 88.92 | 61.70 / 86.53 | 69.26 / 83.45 | 76.07 / 78.82 | 64.37 / 84.43 |
| OptFS | 34.39 / 89.99 | 46.41 / 88.48 | 42.20 / 88.23 | 61.44 / 82.69 | 46.11 / 87.35 |
| AdaSCALE-A | 36.38 / 89.60 | 51.13 / 87.16 | 43.02 / 88.07 | 59.97 / 82.48 | 47.63 / 86.83 |
| AdaSCALE-L | 35.16 / 89.84 | 50.91 / 87.37 | 43.01 / 88.13 | 60.05 / 82.55 | 47.28 / 86.97 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 55.63 / 86.47 | 79.28 / 80.12 | 81.22 / 81.72 | 77.41 / 79.78 | 73.39 / 82.02 |
| MLS | 93.46 / 78.87 | 94.60 / 74.73 | 97.61 / 70.72 | 94.97 / 69.17 | 95.16 / 73.37 |
| EBO | 95.11 / 67.72 | 95.36 / 69.69 | 97.97 / 60.19 | 95.87 / 58.35 | 96.08 / 63.99 |
| ReAct | 40.77 / 88.60 | 62.26 / 85.54 | 58.19 / 85.76 | 74.21 / 79.16 | 58.86 / 84.77 |
| ASH | 98.59 / 42.18 | 98.55 / 43.37 | 98.23 / 43.28 | 97.57 / 43.98 | 98.23 / 43.20 |
| SCALE | 87.83 / 62.98 | 87.71 / 69.63 | 88.75 / 66.63 | 82.08 / 67.82 | 86.59 / 66.77 |
| BFAct | 25.76 / 91.42 | 45.73 / 87.34 | 32.13 / 91.02 | 52.33 / 84.08 | 38.99 / 88.47 |
| LTS | 57.92 / 86.10 | 77.66 / 78.02 | 73.20 / 80.16 | 82.69 / 72.71 | 72.86 / 79.25 |
| OptFS | 31.94 / 90.56 | 50.27 / 86.91 | 36.50 / 90.18 | 58.38 / 83.51 | 44.27 / 87.79 |
| AdaSCALE-A | 32.82 / 90.73 | 61.82 / 85.34 | 38.58 / 89.78 | 58.02 / 82.71 | 47.81 / 87.14 |
| AdaSCALE-L | 30.95 / 91.69 | 60.17 / 86.30 | 37.52 / 90.08 | 56.32 / 83.82 | 46.24 / 87.97 |
E.3 Full-Spectrum near-OOD detection
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 88.17 / 47.34 | 78.15 / 54.73 | 96.29 / 13.81 | 87.54 / 38.63 |
| MLS | 90.04 / 43.32 | 82.06 / 50.23 | 95.59 / 18.94 | 89.23 / 37.50 |
| EBO | 90.19 / 42.62 | 82.64 / 49.01 | 95.54 / 19.57 | 89.46 / 37.07 |
| ReAct | 90.65 / 45.19 | 80.05 / 53.37 | 93.62 / 26.15 | 88.10 / 41.57 |
| ASH | 88.82 / 44.08 | 78.35 / 54.54 | 92.48 / 30.49 | 86.55 / 43.04 |
| SCALE | 85.85 / 48.10 | 77.54 / 57.01 | 93.26 / 32.58 | 85.55 / 45.90 |
| BFAct | 90.43 / 45.29 | 79.62 / 53.50 | 93.62 / 26.20 | 87.89 / 41.66 |
| LTS | 86.37 / 47.43 | 77.54 / 56.40 | 93.61 / 30.57 | 85.84 / 44.80 |
| OptFS | 90.78 / 44.01 | 77.24 / 54.91 | 90.91 / 32.26 | 86.31 / 43.73 |
| AdaSCALE-A | 81.30 / 51.88 | 74.13 / 58.55 | 89.15 / 37.62 | 81.52 / 49.35 |
| AdaSCALE-L | 81.85 / 51.38 | 74.42 / 58.23 | 90.07 / 36.91 | 82.11 / 48.84 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 87.09 / 49.18 | 76.45 / 56.92 | 94.24 / 28.33 | 85.93 / 44.81 |
| MLS | 88.90 / 46.45 | 79.19 / 53.62 | 93.94 / 31.76 | 87.34 / 43.94 |
| EBO | 89.02 / 45.99 | 79.60 / 52.66 | 93.87 / 32.40 | 87.50 / 43.68 |
| ReAct | 89.50 / 47.79 | 77.22 / 56.02 | 89.37 / 39.62 | 85.36 / 47.81 |
| ASH | 87.84 / 46.39 | 75.36 / 56.72 | 88.48 / 42.80 | 83.90 / 48.64 |
| SCALE | 85.81 / 48.94 | 75.33 / 58.79 | 88.49 / 44.45 | 83.21 / 50.73 |
| BFAct | 89.19 / 48.07 | 76.94 / 56.02 | 89.54 / 39.65 | 85.22 / 47.91 |
| LTS | 86.02 / 48.72 | 74.89 / 58.41 | 89.17 / 43.02 | 83.36 / 50.05 |
| OptFS | 89.63 / 46.23 | 75.52 / 56.71 | 85.82 / 44.21 | 83.65 / 49.05 |
| AdaSCALE-A | 82.33 / 51.47 | 74.38 / 59.17 | 82.89 / 48.49 | 79.87 / 53.04 |
| AdaSCALE-L | 82.53 / 51.31 | 75.22 / 58.62 | 82.19 / 48.03 | 79.98 / 52.66 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 83.74 / 57.23 | 72.32 / 67.69 | 87.81 / 56.61 | 81.29 / 60.51 |
| MLS | 82.91 / 60.89 | 71.22 / 70.86 | 93.27 / 55.21 | 82.46 / 62.32 |
| EBO | 82.77 / 61.63 | 71.17 / 71.23 | 93.39 / 55.02 | 82.44 / 62.63 |
| ReAct | 87.74 / 55.64 | 80.22 / 65.24 | 91.11 / 55.13 | 86.36 / 58.67 |
| ASH | 87.26 / 57.45 | 84.81 / 61.59 | 93.78 / 54.76 | 88.61 / 57.93 |
| SCALE | 83.23 / 61.02 | 72.48 / 71.33 | 92.82 / 57.16 | 82.84 / 63.17 |
| BFAct | 90.46 / 54.73 | 87.03 / 61.98 | 92.37 / 54.13 | 89.96 / 56.95 |
| LTS | 83.53 / 60.77 | 74.23 / 70.84 | 92.30 / 57.71 | 83.35 / 63.11 |
| OptFS | 90.33 / 51.78 | 80.82 / 63.86 | 88.86 / 59.45 | 86.67 / 58.36 |
| AdaSCALE-A | 81.68 / 60.46 | 68.05 / 70.95 | 83.25 / 60.91 | 77.66 / 64.11 |
| AdaSCALE-L | 84.30 / 59.42 | 76.30 / 68.49 | 81.86 / 63.12 | 80.82 / 63.68 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 86.95 / 49.77 | 78.27 / 57.25 | 94.79 / 28.95 | 86.67 / 45.32 |
| MLS | 88.56 / 47.94 | 81.58 / 54.31 | 94.26 / 32.40 | 88.13 / 44.88 |
| EBO | 88.68 / 47.69 | 81.67 / 53.51 | 94.21 / 32.99 | 88.19 / 44.73 |
| ReAct | 89.45 / 47.44 | 80.37 / 55.29 | 91.40 / 37.58 | 87.07 / 46.77 |
| ASH | 86.26 / 49.73 | 79.62 / 57.17 | 92.63 / 39.67 | 86.17 / 48.86 |
| SCALE | 84.64 / 52.60 | 78.75 / 58.90 | 94.20 / 37.99 | 85.86 / 49.83 |
| BFAct | 89.22 / 47.70 | 80.39 / 55.34 | 91.24 / 37.66 | 86.95 / 46.90 |
| LTS | 85.03 / 51.87 | 78.65 / 58.48 | 93.77 / 37.96 | 85.82 / 49.43 |
| OptFS | 89.63 / 46.23 | 75.52 / 56.71 | 85.82 / 44.21 | 83.65 / 49.05 |
| AdaSCALE-A | 82.33 / 51.47 | 74.38 / 59.17 | 82.89 / 48.49 | 79.87 / 53.04 |
| AdaSCALE-L | 82.70 / 52.14 | 75.85 / 58.05 | 89.55 / 43.81 | 82.70 / 51.33 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 87.27 / 49.71 | 76.81 / 58.30 | 95.00 / 29.36 | 86.36 / 45.79 |
| MLS | 89.24 / 47.21 | 80.48 / 55.21 | 95.61 / 30.74 | 88.44 / 44.39 |
| EBO | 89.39 / 46.79 | 80.94 / 54.13 | 95.60 / 31.44 | 88.65 / 44.12 |
| ReAct | 90.76 / 45.54 | 79.54 / 55.57 | 88.40 / 42.63 | 86.23 / 47.91 |
| ASH | 89.75 / 47.08 | 81.35 / 58.13 | 90.47 / 46.81 | 87.19 / 50.67 |
| SCALE | 87.55 / 49.53 | 78.35 / 59.54 | 92.79 / 39.22 | 86.23 / 49.43 |
| BFAct | 92.33 / 44.88 | 83.90 / 54.83 | 84.88 / 49.50 | 87.04 / 49.74 |
| LTS | 87.30 / 50.02 | 78.41 / 60.35 | 92.12 / 42.25 | 85.94 / 50.88 |
| OptFS | 92.32 / 43.61 | 81.45 / 56.09 | 85.02 / 50.48 | 86.26 / 50.06 |
| AdaSCALE-A | 86.22 / 48.41 | 80.25 / 56.38 | 81.35 / 47.45 | 82.60 / 50.75 |
| AdaSCALE-L | 86.43 / 48.54 | 80.83 / 56.30 | 81.00 / 50.00 | 82.75 / 51.61 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 83.74 / 57.23 | 72.32 / 67.69 | 87.81 / 56.61 | 81.29 / 60.51 |
| MLS | 82.91 / 60.89 | 71.22 / 70.86 | 93.27 / 55.21 | 82.46 / 62.32 |
| EBO | 82.77 / 61.63 | 71.17 / 71.23 | 93.39 / 55.02 | 82.44 / 62.63 |
| ReAct | 87.74 / 55.64 | 80.22 / 65.24 | 91.11 / 55.13 | 86.36 / 58.67 |
| ASH | 87.26 / 57.45 | 84.81 / 61.59 | 93.78 / 54.76 | 88.61 / 57.93 |
| SCALE | 83.23 / 61.02 | 72.48 / 71.33 | 92.82 / 57.16 | 82.84 / 63.17 |
| BFAct | 90.46 / 54.73 | 87.03 / 61.98 | 92.37 / 54.13 | 89.96 / 56.95 |
| LTS | 83.53 / 60.77 | 74.23 / 70.84 | 92.30 / 57.71 | 83.35 / 63.11 |
| OptFS | 90.33 / 51.78 | 80.82 / 63.86 | 88.86 / 59.45 | 86.67 / 58.36 |
| AdaSCALE-A | 81.68 / 60.46 | 68.05 / 70.95 | 83.25 / 60.91 | 77.66 / 64.11 |
| AdaSCALE-L | 84.30 / 59.42 | 76.30 / 68.49 | 81.86 / 63.12 | 80.82 / 63.68 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 92.28 / 47.57 | 87.44 / 56.23 | 98.02 / 39.33 | 92.58 / 47.71 |
| MLS | 94.11 / 44.88 | 95.17 / 52.44 | 98.00 / 37.77 | 95.76 / 45.03 |
| EBO | 94.47 / 42.06 | 95.86 / 48.45 | 97.89 / 38.03 | 96.07 / 42.85 |
| ReAct | 94.95 / 41.84 | 88.65 / 52.48 | 95.21 / 44.64 | 92.94 / 46.32 |
| ASH | 88.95 / 56.47 | 91.09 / 55.11 | 90.00 / 55.78 | 90.01 / 55.79 |
| SCALE | 94.52 / 41.21 | 96.30 / 45.78 | 97.76 / 37.09 | 96.19 / 41.36 |
| BFAct | 94.99 / 41.44 | 85.62 / 53.35 | 92.62 / 45.64 | 91.07 / 46.81 |
| LTS | 95.33 / 43.36 | 90.52 / 53.14 | 95.90 / 41.30 | 93.91 / 45.93 |
| OptFS | 94.19 / 43.01 | 81.66 / 55.60 | 88.90 / 46.04 | 88.25 / 48.22 |
| AdaSCALE-A | 93.30 / 42.49 | 81.00 / 54.72 | 84.12 / 45.71 | 86.14 / 47.64 |
| AdaSCALE-L | 93.52 / 41.83 | 80.94 / 54.16 | 84.26 / 45.82 | 86.24 / 47.27 |
| Method | SSB-Hard | NINCO | ImageNet-O | Average |
| MSP | 91.55 / 53.29 | 86.73 / 60.62 | 97.85 / 42.90 | 92.04 / 52.27 |
| MLS | 94.49 / 50.01 | 93.94 / 56.40 | 97.11 / 43.76 | 95.18 / 50.06 |
| EBO | 94.66 / 47.41 | 94.58 / 52.04 | 96.76 / 45.84 | 95.33 / 48.43 |
| ReAct | 94.04 / 47.83 | 82.85 / 58.52 | 94.60 / 50.41 | 90.50 / 52.25 |
| ASH | 91.77 / 50.35 | 90.91 / 52.09 | 88.80 / 54.50 | 90.49 / 52.31 |
| SCALE | 93.39 / 47.38 | 91.25 / 53.00 | 90.80 / 56.10 | 91.81 / 52.16 |
| BFAct | 92.65 / 48.33 | 79.61 / 59.66 | 84.26 / 53.17 | 85.51 / 53.72 |
| LTS | 94.26 / 48.70 | 88.32 / 57.60 | 93.04 / 46.33 | 91.87 / 50.88 |
| OptFS | 94.06 / 47.77 | 81.91 / 59.14 | 86.95 / 51.34 | 87.64 / 52.75 |
| AdaSCALE-A | 90.29 / 46.84 | 81.54 / 57.44 | 87.74 / 47.55 | 86.52 / 50.61 |
| AdaSCALE-L | 90.18 / 46.63 | 80.77 / 57.85 | 87.27 / 47.90 | 86.07 / 50.79 |
E.4 Full-Spectrum far-OOD detection
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 69.31 / 65.65 | 80.57 / 59.22 | 73.94 / 60.74 | 79.02 / 56.04 | 75.71 / 60.41 |
| MLS | 64.71 / 63.30 | 74.69 / 61.67 | 69.73 / 60.60 | 80.04 / 55.08 | 72.29 / 60.16 |
| EBO | 65.30 / 61.43 | 74.48 / 61.87 | 69.92 / 59.93 | 80.12 / 54.48 | 72.45 / 59.42 |
| ReAct | 51.90 / 75.79 | 63.55 / 69.22 | 65.64 / 67.36 | 71.81 / 67.01 | 63.22 / 69.84 |
| ASH | 49.21 / 76.93 | 50.54 / 77.64 | 63.04 / 69.03 | 70.62 / 64.79 | 58.35 / 72.10 |
| SCALE | 43.34 / 79.23 | 46.60 / 79.58 | 62.26 / 70.54 | 67.94 / 67.05 | 55.04 / 74.10 |
| BFAct | 51.01 / 75.85 | 62.45 / 69.03 | 65.52 / 67.21 | 71.17 / 66.45 | 62.54 / 69.63 |
| LTS | 45.12 / 78.72 | 48.77 / 79.13 | 62.43 / 70.17 | 69.31 / 66.20 | 56.41 / 73.55 |
| OptFS | 49.39 / 77.14 | 50.25 / 75.59 | 62.46 / 68.87 | 69.84 / 65.65 | 57.99 / 71.81 |
| AdaSCALE-A | 41.24 / 79.53 | 45.55 / 79.64 | 56.43 / 72.85 | 66.13 / 67.54 | 52.33 / 74.89 |
| AdaSCALE-L | 41.75 / 79.63 | 45.67 / 79.96 | 56.80 / 72.82 | 66.69 / 67.28 | 52.73 / 74.92 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 71.78 / 64.51 | 78.82 / 62.20 | 72.52 / 62.27 | 78.71 / 57.55 | 75.46 / 61.63 |
| MLS | 70.55 / 62.07 | 72.11 / 65.46 | 68.68 / 62.27 | 77.60 / 58.03 | 72.23 / 61.96 |
| EBO | 70.94 / 60.64 | 72.18 / 65.71 | 68.96 / 61.59 | 77.97 / 57.75 | 72.51 / 61.42 |
| ReAct | 53.59 / 75.05 | 59.92 / 72.13 | 62.45 / 69.16 | 71.14 / 66.87 | 61.78 / 70.80 |
| ASH | 53.84 / 74.07 | 47.67 / 79.09 | 60.66 / 70.12 | 71.45 / 64.40 | 58.40 / 71.92 |
| SCALE | 47.82 / 76.79 | 42.04 / 80.95 | 59.26 / 71.76 | 70.32 / 65.86 | 54.86 / 73.84 |
| BFAct | 53.23 / 74.86 | 58.84 / 71.86 | 62.33 / 68.97 | 70.67 / 66.17 | 61.27 / 70.47 |
| LTS | 49.64 / 76.17 | 43.87 / 80.60 | 59.32 / 71.46 | 70.85 / 65.35 | 55.92 / 73.39 |
| OptFS | 51.52 / 75.52 | 48.88 / 76.96 | 59.99 / 70.28 | 70.86 / 65.15 | 57.81 / 71.98 |
| AdaSCALE-A | 44.77 / 77.51 | 42.17 / 80.73 | 53.58 / 73.90 | 67.14 / 66.84 | 51.91 / 74.75 |
| AdaSCALE-L | 46.52 / 76.39 | 44.87 / 79.90 | 53.32 / 73.81 | 68.16 / 65.88 | 53.22 / 73.99 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 53.98 / 80.49 | 69.36 / 70.97 | 62.02 / 75.95 | 75.48 / 65.76 | 65.21 / 73.29 |
| MLS | 39.99 / 85.09 | 69.12 / 73.89 | 58.30 / 80.08 | 79.98 / 66.96 | 61.85 / 76.51 |
| EBO | 38.42 / 86.11 | 68.23 / 74.16 | 58.77 / 80.93 | 80.56 / 67.08 | 61.49 / 77.07 |
| ReAct | 43.95 / 85.15 | 66.24 / 73.34 | 68.23 / 77.80 | 88.79 / 57.47 | 66.80 / 73.44 |
| ASH | 61.03 / 79.08 | 58.16 / 80.78 | 80.03 / 74.57 | 81.60 / 67.74 | 70.21 / 75.54 |
| SCALE | 39.13 / 86.34 | 56.30 / 79.93 | 60.66 / 81.10 | 76.25 / 70.04 | 58.09 / 79.35 |
| BFAct | 52.17 / 82.68 | 71.55 / 70.74 | 78.37 / 74.00 | 90.50 / 56.09 | 73.15 / 70.88 |
| LTS | 38.98 / 86.90 | 50.10 / 82.26 | 65.29 / 81.17 | 75.41 / 70.91 | 57.44 / 80.31 |
| OptFS | 52.61 / 82.40 | 62.47 / 76.57 | 66.68 / 76.79 | 88.01 / 56.14 | 67.44 / 72.98 |
| AdaSCALE-A | 34.25 / 87.68 | 63.81 / 76.12 | 50.37 / 82.13 | 74.81 / 68.37 | 55.81 / 78.58 |
| AdaSCALE-L | 32.72 / 88.82 | 47.93 / 82.84 | 53.84 / 83.09 | 73.90 / 70.37 | 52.10 / 81.28 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 68.90 / 66.67 | 80.91 / 60.21 | 71.91 / 63.25 | 78.58 / 57.84 | 75.07 / 62.00 |
| MLS | 64.59 / 65.49 | 76.51 / 62.51 | 67.70 / 63.96 | 79.89 / 56.90 | 72.17 / 62.21 |
| EBO | 64.95 / 64.23 | 76.59 / 62.61 | 68.00 / 63.52 | 79.80 / 56.39 | 72.34 / 61.69 |
| ReAct | 51.97 / 76.08 | 65.34 / 68.27 | 63.05 / 68.94 | 70.13 / 67.58 | 62.62 / 70.22 |
| ASH | 53.84 / 74.07 | 47.67 / 79.09 | 60.66 / 70.12 | 71.45 / 64.40 | 58.40 / 71.92 |
| SCALE | 49.27 / 76.82 | 60.07 / 75.14 | 62.91 / 70.69 | 73.38 / 64.22 | 61.41 / 71.72 |
| BFAct | 53.23 / 74.86 | 58.84 / 71.86 | 62.33 / 68.97 | 70.67 / 66.17 | 61.27 / 70.47 |
| LTS | 49.64 / 76.17 | 48.77 / 79.13 | 62.43 / 70.17 | 70.85 / 65.35 | 56.41 / 73.55 |
| OptFS | 51.52 / 75.77 | 50.25 / 75.59 | 62.46 / 68.87 | 69.84 / 65.65 | 57.99 / 71.81 |
| AdaSCALE-A | 43.82 / 78.47 | 45.55 / 79.64 | 56.43 / 72.85 | 66.13 / 67.54 | 52.33 / 74.89 |
| AdaSCALE-L | 46.23 / 76.67 | 54.29 / 75.80 | 56.84 / 72.44 | 69.13 / 64.99 | 56.62 / 72.47 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 68.90 / 66.67 | 80.91 / 60.21 | 71.91 / 63.25 | 78.58 / 57.84 | 75.07 / 62.00 |
| MLS | 64.59 / 65.49 | 76.51 / 62.51 | 67.70 / 63.96 | 79.89 / 56.90 | 72.17 / 62.21 |
| EBO | 64.95 / 64.23 | 76.59 / 62.61 | 68.00 / 63.52 | 79.80 / 56.39 | 72.34 / 61.69 |
| ReAct | 51.97 / 76.08 | 65.34 / 68.27 | 63.05 / 68.94 | 70.13 / 67.58 | 62.62 / 70.22 |
| ASH | 53.84 / 74.07 | 47.67 / 79.09 | 60.66 / 70.12 | 71.45 / 64.40 | 58.40 / 71.92 |
| SCALE | 49.27 / 76.82 | 60.07 / 75.14 | 62.91 / 70.69 | 73.38 / 64.22 | 61.41 / 71.72 |
| BFAct | 53.23 / 74.86 | 58.84 / 71.86 | 62.33 / 68.97 | 70.67 / 66.17 | 61.27 / 70.47 |
| LTS | 49.64 / 76.17 | 48.77 / 79.13 | 62.43 / 70.17 | 70.85 / 65.35 | 56.41 / 73.55 |
| OptFS | 51.52 / 75.77 | 50.25 / 75.59 | 62.46 / 68.87 | 69.84 / 65.65 | 57.99 / 71.81 |
| AdaSCALE-A | 52.55 / 73.62 | 54.70 / 76.19 | 58.11 / 71.31 | 77.77 / 58.83 | 60.78 / 69.99 |
| AdaSCALE-L | 53.04 / 73.62 | 51.83 / 78.00 | 58.30 / 71.92 | 78.40 / 58.47 | 60.39 / 70.50 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 53.98 / 80.49 | 69.36 / 70.97 | 62.02 / 75.95 | 75.48 / 65.76 | 65.21 / 73.29 |
| MLS | 39.99 / 85.09 | 69.12 / 73.89 | 58.30 / 80.08 | 79.98 / 66.96 | 61.85 / 76.51 |
| EBO | 38.42 / 86.11 | 68.23 / 74.16 | 58.77 / 80.93 | 80.56 / 67.08 | 61.49 / 77.07 |
| ReAct | 43.95 / 85.15 | 66.24 / 73.34 | 68.23 / 77.80 | 88.79 / 57.47 | 66.80 / 73.44 |
| ASH | 61.03 / 79.08 | 58.16 / 80.78 | 80.03 / 74.57 | 81.60 / 67.74 | 70.21 / 75.54 |
| SCALE | 39.13 / 86.34 | 56.30 / 79.93 | 60.66 / 81.10 | 76.25 / 70.04 | 58.09 / 79.35 |
| BFAct | 52.17 / 82.68 | 71.55 / 70.74 | 78.37 / 74.00 | 90.50 / 56.09 | 73.15 / 70.88 |
| LTS | 38.98 / 86.90 | 50.10 / 82.26 | 65.29 / 81.17 | 75.41 / 70.91 | 57.44 / 80.31 |
| OptFS | 52.61 / 82.40 | 62.47 / 76.57 | 66.68 / 76.79 | 88.01 / 56.14 | 67.44 / 72.98 |
| AdaSCALE-A | 40.76 / 88.62 | 63.95 / 77.69 | 53.96 / 84.71 | 77.17 / 69.22 | 58.96 / 80.06 |
| AdaSCALE-L | 43.28 / 87.91 | 50.23 / 83.05 | 57.11 / 84.45 | 73.87 / 70.29 | 56.12 / 81.43 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 66.12 / 67.29 | 75.49 / 64.02 | 75.32 / 63.46 | 83.84 / 58.77 | 75.19 / 63.39 |
| MLS | 82.34 / 64.58 | 85.83 / 63.52 | 90.12 / 61.11 | 92.95 / 55.08 | 87.81 / 61.07 |
| EBO | 87.94 / 59.51 | 88.03 / 62.20 | 91.84 / 57.54 | 94.14 / 50.68 | 90.49 / 57.48 |
| ReAct | 70.31 / 62.56 | 75.49 / 64.90 | 76.64 / 61.30 | 87.02 / 54.77 | 77.37 / 60.88 |
| ASH | 93.23 / 53.23 | 95.19 / 51.16 | 90.38 / 57.95 | 89.04 / 56.56 | 91.96 / 54.72 |
| SCALE | 90.13 / 55.74 | 88.71 / 61.09 | 92.30 / 55.21 | 94.75 / 47.62 | 91.47 / 54.92 |
| BFAct | 66.39 / 62.89 | 71.84 / 65.55 | 71.56 / 62.19 | 84.23 / 55.84 | 73.51 / 61.62 |
| LTS | 71.01 / 67.15 | 78.15 / 65.01 | 82.72 / 60.98 | 86.73 / 56.56 | 79.65 / 62.43 |
| OptFS | 62.89 / 66.41 | 70.96 / 65.68 | 68.25 / 64.30 | 80.02 / 58.53 | 70.53 / 63.73 |
| AdaSCALE-A | 65.49 / 65.27 | 74.75 / 63.69 | 69.79 / 63.49 | 79.93 / 57.47 | 72.49 / 62.48 |
| AdaSCALE-L | 64.72 / 64.84 | 74.72 / 63.49 | 69.92 / 62.86 | 79.99 / 57.04 | 72.34 / 62.06 |
| Method | iNaturalist | Textures | OpenImage-O | Places | Average |
| MSP | 73.78 / 70.69 | 87.48 / 63.62 | 88.58 / 65.05 | 86.43 / 62.22 | 84.07 / 65.39 |
| MLS | 94.01 / 64.59 | 94.98 / 61.02 | 97.72 / 57.01 | 95.29 / 54.55 | 95.50 / 59.29 |
| EBO | 95.11 / 55.73 | 95.35 / 58.70 | 97.99 / 49.71 | 95.87 / 47.30 | 96.08 / 52.86 |
| ReAct | 66.17 / 68.53 | 79.27 / 66.85 | 76.92 / 65.82 | 85.99 / 57.89 | 77.09 / 64.77 |
| ASH | 94.55 / 47.18 | 94.43 / 48.22 | 93.77 / 48.28 | 92.55 / 48.99 | 93.82 / 48.17 |
| SCALE | 91.23 / 53.32 | 91.15 / 60.76 | 91.87 / 57.44 | 87.05 / 58.13 | 90.32 / 57.41 |
| BFAct | 54.53 / 73.59 | 69.73 / 68.65 | 59.76 / 74.29 | 74.09 / 64.30 | 64.53 / 70.21 |
| LTS | 72.97 / 70.44 | 86.21 / 62.18 | 83.29 / 63.87 | 89.39 / 56.32 | 82.96 / 63.20 |
| OptFS | 58.91 / 71.84 | 72.14 / 68.00 | 62.45 / 72.16 | 77.16 / 63.35 | 67.66 / 68.83 |
| AdaSCALE-A | 61.33 / 68.58 | 79.90 / 64.03 | 65.40 / 67.52 | 77.68 / 59.21 | 71.08 / 64.83 |
| AdaSCALE-L | 60.15 / 70.58 | 78.72 / 65.23 | 64.83 / 68.17 | 76.43 / 60.64 | 70.03 / 66.15 |

### v2

@@
 These approaches resort to sample-specific scaling but apply a static percentile threshold across all samples regardless of their nature, resulting in suboptimal ID-OOD separability.
-In this work, we propose AdaSCALE, an adaptive scaling procedure that dynamically adjusts the percentile threshold based on a sample’s estimated OOD likelihood.
+In this work, we propose AdaSCALE, an adaptive scaling procedure that dynamically adjusts the percentile threshold based on a sample’s estimated OODness.
 This estimation leverages our key observation: OOD samples exhibit significantly more pronounced activation shifts at high-magnitude activations under minor perturbation compared to ID samples.
@@
 The reliable deployment of deep learning models hinges on their ability to handle previously unseen inputs, a task commonly known as OOD detection.
-One critical application is in medical diagnosis, where a model trained on common diseases should be able to flag inputs representing unknown conditions as potential outliers, requiring further review by clinicians.
-OOD detection primarily involves identifying semantic shifts, with robustness to covariate shifts being a highly desirable characteristic [58, 2].
-As modern deep learning models scale in both data and parameter counts, effective OOD detection within large-scale settings is critical.
-Given the difficulties of iterating on large models, post-hoc approaches that preserve ID accuracy are generally preferred.
+A critical application is in medical diagnosis, where a model trained on common diseases should be able to flag inputs representing unknown conditions as potential outliers, requiring further review by clinicians.
+OOD detection primarily involves identifying semantic shifts, with robustness to covariate shifts being a highly desirable characteristic (Yang et al., 2023; Baek et al., 2024).
+As deep learning models scale in both data and parameter counts, effective OOD detection within large-scale settings is critical.
+Given the difficulties of iterating on large models, post-hoc approaches preserving ID accuracy are generally preferred.
 A variety of post-hoc approaches have emerged, broadly categorized by where they operate.
-One class of methods focuses on computing OOD scores directly in the output space [16, 30, 33, 20, 7, 64], while another operates in the activation space [27, 48, 44, 40].
-Finally, a more recent line of research also explores a hybrid approach [50, 24], combining information from both spaces.
-The efficacy of many high-performing methods relies on either accurate computation of ID statistics [47, 25, 55, 46, 26] or retention of training data statistics [48, 40].
-However, as retaining full access to training data becomes increasingly impractical in large-scale settings, methods that operate effectively with minimal ID samples without performance degradation are particularly valuable for practical applications.
-Alleviating the dependence on ID training data/statistics, recent state-of-the-art post-hoc approaches center around the concept of “fixed scaling.” ASH [7] prunes and scales activations on a per-sample basis.
-SCALE [54], the direct successor of ASH, critiques pruning and focuses purely on scaling, which improves OOD detection without accuracy degradation.
-LTS [8] extends this concept by directly scaling logits instead, using post-ReLU activations.
-These methods leverage a key insight: scaling based on the relative strength of a sample’s top-k activations (with respect to the entire activations) produces highly separable ID-OOD energy scores.
-However, although such approaches provide sample-specific scaling factors, the scaling mechanism remains uniform across all samples as the percentile threshold and thereby is fixed, as shown in Figure 1.
+One class of methods focuses on computing OOD scores directly in the output space (Hendrycks & Gimpel, 2017; Liang et al., 2018; Liu et al., 2020; Hendrycks et al., 2022a; Djurisic et al., 2023; Zhao et al., 2024), while another operates in the activation space (Lee et al., 2018; Sun et al., 2022; Ren et al., 2021; Rajasekaran et al., 2024).
+Finally, a more recent line of research also explores a hybrid approach (Wang et al., 2022; Kim et al., 2024), combining information from both spaces.
+The efficacy of many high-performing methods relies on either accurate computation of ID statistics (Sun et al., 2021; Kong & Li, 2022; Xu et al., 2023; Sun & Li, 2022; Krumpl et al., 2024) or retention of training data statistics (Sun et al., 2022; Rajasekaran et al., 2024).
+However, as retaining full access to training data becomes increasingly impractical in large-scale settings, methods that operate effectively with minimal ID samples are particularly valuable for practical applications.
+Alleviating the dependence on ID training data/statistics, recent state-of-the-art post-hoc approaches center around the concept of “fixed scaling." ASH (Djurisic et al., 2023) prunes and scales activations on a per-sample basis.
+SCALE (Xu et al., 2024), the direct successor of ASH, critiques pruning and focuses purely on scaling, which improves OOD detection performance without accuracy degradation.
+LTS (Djurisic et al., 2024) extends this concept by directly scaling logits instead, using post-ReLU activations.
+These methods leverage a key insight: scaling based on relative strength of a sample’s top-k activations (with respect to entire activations) yields highly separable ID-OOD energy scores.
+However, although such approaches provide sample-specific scaling factors, the scaling mechanism remains uniform across all samples as the percentile threshold and thereby is fixed, as shown in Figure˜1.
 This static approach is inherently limiting for optimal ID-OOD separation while also failing to leverage even minimal ID data, which could be reasonably practical in most deployment scenarios.
-We hypothesize that designing an adaptive scaling procedure based on each sample’s predetermined OOD likelihood offers greater control for enhancing ID-OOD separability.
-Specifically, this mechanism should assign smaller scaling factors for samples with high OOD likelihood to yield lower energy scores and larger scaling factors for probable ID samples to yield higher energy scores.
-To achieve this, we propose a heuristic for predetermining OOD likelihood based on a key observation in activation space: minor perturbations applied to OOD samples induce significantly more pronounced shifts in their top-k activations compared to ID samples.
+We hypothesize that designing an adaptive scaling procedure based on each sample’s estimated OODness offers greater control for enhancing ID-OOD separability.
+Specifically, this mechanism should assign smaller scaling factors for samples with high OODness to yield lower energy scores and larger scaling factors for probable ID samples to yield higher energy scores.
+To achieve this, we propose a heuristic for estimating OODness based on a key observation in activation space: minor perturbations applied to OOD samples induce significantly more pronounced shifts in their top-k activations compared to ID samples.
 Consequently, samples exhibiting substantial activation shifts are assigned lower scaling factors, while those with minimal shifts receive higher scaling factors.
@@
 We conduct an extensive evaluation across 8 architectures on ImageNet-1k and 2 architectures on CIFAR benchmarks, demonstrating the substantial effectiveness of AdaSCALE.
-For instance, AdaSCALE surpasses the average performance of the best-generalizing method, OptFS [64], by for near-OOD detection and for far-OOD detection in terms of FPR@95 / AUROC, on the ImageNet-1k benchmark across eight architectures.
-Furthermore, AdaSCALE outperforms the best-performing method, SCALE [54], when evaluated on the ResNet-50 architecture, achieving performance gains of for near-OOD and for far-OOD detection.
-Additionally, AdaSCALE consistently demonstrates superiority in full-spectrum OOD detection [58].
-Our key contributions are summarized as follows:
-We reveal that OOD inputs exhibit significantly more pronounced shifts in top-k activations under minor perturbations compared to ID inputs.
-Leveraging this observation, we propose a novel post-hoc OOD detection method utilizing adaptive scaling that attains state-of-the-art OOD detection.
-We demonstrate state-of-the-art generalization of AdaSCALE via extensive evaluations across many setups.
+For instance, AdaSCALE surpasses the average performance of the best-generalizing method, OptFS (Zhao et al., 2024), by for near-OOD detection and for far-OOD detection in terms of FPR@95 / AUROC, on the ImageNet-1k benchmark across eight architectures.
+Furthermore, AdaSCALE outperforms the best-performing method, SCALE (Xu et al., 2024), when evaluated on the ResNet-50 architecture, achieving performance gains of for near-OOD and for far-OOD detection.
+Additionally, AdaSCALE consistently demonstrates superiority in full-spectrum OOD (FSOOD) detection (Yang et al., 2023).
+Our key contributions are:
+We reveal that OOD inputs exhibit more pronounced shifts in top-k activations under minor perturbations compared to ID inputs.
+Leveraging this, we propose a novel post-hoc OOD detection method using adaptive scaling that attains state-of-the-art OOD detection.
+We demonstrate state-of-the-art generalization of AdaSCALE via extensive evaluations across 10 architectures and 3 datasets by tuning mere one hyperparameter for a given setup.
 2 Related Works
 Post-hoc methods.
-Early research on OOD detection primarily focused on designing scoring functions based on logit information [16, 30, 33, 20, 34].
-While these methods leveraged logit-based scores, alternative approaches have explored gradient-based information, such as GradNorm [23], GradOrth [4], GAIA [5], and Greg-OOD [45].
+Early research on OOD detection primarily focused on designing scoring functions based on logit information (Hendrycks & Gimpel, 2017; Liang et al., 2018; Liu et al., 2020; Hendrycks et al., 2022a; Liu et al., 2023).
+While these methods leveraged logit-based scores, alternative approaches have explored gradient-based information, such as GradNorm (Huang et al., 2021), GradOrth (Behpour et al., 2023), GAIA (Chen et al., 2023), and Greg-OOD (Sharifi et al., 2025).
 Given the limited dimensionality of the logit space, which may not encapsulate sufficient information for OOD detection, subsequent studies have investigated activation-space-based methods.
-These approaches exploit the high-dimensional activations, leading to both parametric techniques such as MDS [27], MDS Ensemble [27], and RMDS [44], as well as non-parametric methods such as KNN-based OOD detection [48, 39].
+These approaches exploit the high-dimensional activations, leading to both parametric techniques such as MDS (Lee et al., 2018), MDS Ensemble (Lee et al., 2018), and RMDS (Ren et al., 2021), as well as non-parametric methods such as KNN-based OOD detection (Sun et al., 2022; Park et al., 2023).
 Recent advancements have proposed hybrid methodologies that integrate parametric and non-parametric techniques to improve robustness.
-For instance, ComboOOD [40] combines these paradigms to enhance near-OOD detection performance.
-Similarly, VIM [50] employs a combination of logit-based and distance-based metrics.
-However, reliance of such approaches on ID statistics [46, 38, 59] can become a constraint, hindering scalability and practical deployment in real-world applications.
-To mitigate computational challenges for real-world deployment, recent methods, such as FDBD [32] and NCI [31], have focused on enhancing efficiency.
-Recent advances, such as WeiPer [14], explore class-direction perturbations, while NECO [1] examines connections to neural collapse phenomena.
+For instance, ComboOOD (Rajasekaran et al., 2024) combines these paradigms to enhance near-OOD detection performance.
+Similarly, VIM (Wang et al., 2022) employs a combination of logit-based and distance-based metrics.
+However, reliance of such approaches on ID statistics (Sun & Li, 2022; Olber et al., 2023; Zhang et al., 2023c) can become a constraint, hindering scalability and practical deployment in real-world applications.
+To mitigate computational challenges for real-world deployment, recent methods, such as FDBD (Liu & Qin, 2024) and NCI (Liu & Qin, 2023), have focused on enhancing efficiency.
+Recent advances, such as NECO (Ammar et al., 2024) examines connections to neural collapse phenomena, while WeiPer (Granz et al., 2024), explore class-direction perturbations.
+Unlike WeiPer, our work deals with perturbation in the input image similar to ODIN (Liang et al., 2018).
 Activation-shaping post-hoc methods.
-A seminal work in OOD detection, ReAct [47], identified abnormally high activation patterns in OOD samples and proposed clipping extreme activations.
-This approach has been further generalized by BFAct [25] and VRA [55], which extend activation clipping for enhanced effectiveness.
-Additionally, BATS [66] refines activation distributions by aligning them with their respective typical sets, while LAPS [15] enhances this strategy by incorporating channel-aware typical sets.
-Inspired by activation clipping, another line of research explores activation “scaling” as a means to improve OOD detection.
-ASH [7] introduces a method to compute a scaling factor as a function of the activation itself, pruning and rescaling activations to enhance the separation of energy scores between ID and OOD samples.
+A seminal work in OOD detection, ReAct (Sun et al., 2021), identified abnormally high activation patterns in OOD samples and proposed clipping extreme activations.
+This approach has been further generalized by BFAct (Kong & Li, 2022) and VRA (Xu et al., 2023), which extend activation clipping for enhanced effectiveness.
+Additionally, BATS (Zhu et al., 2022) refines activation distributions by aligning them with their respective typical sets, while LAPS (He et al., 2024) enhances this strategy by incorporating channel-aware typical sets.
+Inspired by activation clipping, another line of research explores activation “scaling" as a means to improve OOD detection.
+ASH (Djurisic et al., 2023) introduces a method to compute a scaling factor as a function of the activation itself, pruning and rescaling activations to enhance the separation of energy scores between ID and OOD samples.
 However, this approach results in a slight degradation in ID classification accuracy.
-In response, SCALE [54] observes that pruning adversely affects performance and thus eliminates it, leading to improved OOD detection while preserving ID accuracy.
+In response, SCALE (Xu et al., 2024) observes that pruning adversely affects performance and thus eliminates it, leading to improved OOD detection while preserving ID accuracy.
 SCALE currently represents the state-of-the-art method for ResNet-50-based OOD detection.
 Despite their efficacy, these activation-based methods exhibit limited generalization across diverse architectures.
-To address this issue, LTS [8] extends SCALE by computing scaling factors using post-ReLU activations and applying them directly to logits rather than activations.
+To address this issue, LTS (Djurisic et al., 2024) extends SCALE by computing scaling factors using post-ReLU activations and applying them directly to logits rather than activations.
 Our work builds on this line of work, introducing the adaptive scaling mechanism.
-ATS [26] argues that relying solely on final-layer activations may result in the loss of critical information beneficial for OOD detection and proposes to leverage intermediate-layer activations too.
+ATS (Krumpl et al., 2024) argues that relying solely on final-layer activations may result in the loss of critical information beneficial for OOD detection and proposes to leverage intermediate-layer activations too.
 However, its efficacy is contingent upon the availability of a large number of training samples, whereas our approach attains state-of-the-art performance while utilizing a minimal number of ID samples.
-A newly proposed method OptFS [64] introduces a piecewise constant shaping function with the goal of generalization across diverse architectures in large-scale settings, while our work exhibits superior generalization extending to small-scale settings too.
+A newly proposed method OptFS (Zhao et al., 2024) introduces a piecewise constant shaping function with the goal of generalization across diverse architectures in large-scale settings, while our work exhibits superior generalization extending to small-scale settings too.
 Training methods.
 The training methods incorporate adjustments during training to enhance the ID-OOD differentiating characteristics.
-They either make architectural adjustments [6, 18, 22], apply enhanced data augmentations [53, 19, 21], or make simple training modifications [52, 42, 62].
-More recent methods have adopted contrastive learning in the context of OOD detection [36, 43, 35, 67].
-Moreover, some approaches also either utilize external real outliers [17, 60, 11, 65] or synthesize virtual outliers either in image space [10, 41, 51, 13, 63, 29, 3, 37] or in feature space [9, 49, 12, 28].
-However, training strategies incur additional computational cost and potentially inferior performance to post-hoc methods in large-scale setups [57].
+They either make architectural adjustments (DeVries & Taylor, 2018; Hendrycks et al., 2019b; Hsu et al., 2020), apply enhanced data augmentations (Xiong et al., 2024; Hendrycks* et al., 2020; Hendrycks et al., 2022b), or make simple training modifications Wei et al.
+(2022); Regmi et al.
+(2024a); Zhang et al.
+(2024).
+More recent methods have adopted contrastive learning in the context of OOD detection (Ming et al., 2023; Regmi et al., 2024b; Lu et al., 2024; Zou et al., 2025).
+Moreover, some approaches also either utilize external real outliers (Hendrycks et al., 2019a; Zhang et al., 2023a; Du et al., 2024; Zhu et al., 2023) or synthesize virtual outliers either in image space (Du et al., 2023; Regmi, 2024; Wang et al., 2023; Gao et al., 2023; Zhang et al., 2025; Li et al., 2024; Bai et al., 2024; Nie et al., 2024) or in feature space (Du et al., 2022; Tao et al., 2023; Gao et al., 2024; Li & Zhang, 2025).
+However, training methods can be costlier and less effective than post-hoc approaches in some large-scale setups (Yang et al., 2022).
 3 Preliminaries
@@
 We refer to individual dimensions of the feature vector as activations, denoted by for the -th dimension.
-The classifier is trained on to minimize the empirical risk:
-where is a loss function, such as cross-entropy loss.
+The classifier is trained on to minimize the empirical risk: where is a loss function, such as cross-entropy loss.
 During inference, the model may encounter data points drawn from a different distribution, denoted as , which is referred to as OOD data.
@@
 A higher score typically indicates a higher probability of the input being OOD.
-A threshold is used to classify an input as either ID or OOD:
+A threshold is used to classify an input as either ID or OOD: .
 4 Method
-In this section, we introduce AdaSCALE, a novel post-processing approach that dynamically adapts the scaling mechanism based on each sample’s estimated OOD likelihood.
-We first present our key empirical observations regarding activation behavior under minor perturbations, building upon insights from ReAct [47].
-Next, we revisit and analyze the core principle underlying recent state-of-the-art approaches.
+In this section, we introduce AdaSCALE, a novel post-processing approach that dynamically adapts the scaling mechanism based on each sample’s estimated OODness.
+We first revisit and analyze the core principle underlying recent scaling-based static state-of-the-art approaches.
+Then, we present our key empirical observations regarding activation behavior under minor perturbations, building upon insights from ReAct (Sun et al., 2021).
 Finally, we detail our proposed adaptive scaling mechanism that leverages these observations to achieve superior OOD detection performance.
+4.1 Revisiting Static Scaling Mechanism
+Scaling baselines (Djurisic et al., 2023; Xu et al., 2024; Djurisic et al., 2024) use energy score on (directly or indirectly) scaled logits, with higher (magnitude) values indicating higher IDness.
+They operate by scaling activations / logits with scaling factor computed as:
+where denotes the percentile of all elements in activation .
+While this approach yields sample-specific scaling factors, it imposes a critical constraint: the percentile threshold is static and identical across all test samples, regardless of the nature of samples.
+We argue that this static nature limits the effectiveness of the scaling procedure and prevents optimal ID-OOD separability.
+We hypothesize that a proxy for estimating the OODness can be useful for computing dynamic percentiles giving rise to dynamic scaling factors.
 4.2 Observations in Activation Space
-A seminal work ReAct [47] demonstrated that OOD samples often induce abnormally high activations within neural networks.
+A seminal work ReAct (Sun et al., 2021) demonstrated that OOD samples often induce abnormally high activations within neural networks.
 We extend this finding with an important observation: the positions of such high activations in OOD samples are relatively unstable under minor perturbations compared to ID samples.
 This instability provides a valuable signal for distinguishing OOD samples from ID samples.
-Below, we formalize this observation and our methodology.
+Below, we formalize this observation.
 4.2.1 Perturbation Mechanism
@@
 We select of channel value indices with lowest absolute attribution scores, forming the set .
-We use a gradient-based attribution:
-where is the predicted class index. To create a perturbed input, we select a subset containing of channel values to perturb. The perturbed image is obtained as:
-where is perturbation magnitude.
+We use a gradient-based attribution , where is predicted class index.
+To create a perturbed input, we select a subset containing of channel values to perturb.
+If is perturbation magnitude, the perturbed image is obtained as:
 4.2.2 Activation Shift as OOD Indicator
-After obtaining the perturbed input , we compute its activation . We define the activation shift as the absolute element-wise difference between the original activation and the perturbed activation:
+After obtaining perturbed input , we compute its activation . We define activation shift as the absolute element-wise difference between original activation and the perturbed one:
 Figure 2 illustrates the key insight of our approach: activation shift at extreme (high-magnitude) activations is consistently more pronounced in OOD samples compared to ID samples.
 This behavior can be understood intuitively: ID samples activate network features in a stable, predictable manner reflecting learned patterns, while OOD samples trigger less stable, more arbitrary high activations that shift significantly under perturbation.
-Based on this observation, we propose using activation shift at the top- highest activations as a metric to estimate OOD likelihood of a sample:
+Based on this observation, we propose using activation shift at the top- highest activations as a metric to estimate OODness of a sample:
 where returns the indices of the highest values in .
-As evidenced by ratio shown in Figure 4, the statistic generally assigns higher values to diverse OOD samples than ID samples.
-However, the high variance of metric (Figure 3) suggests the possibility of overoptimistic estimations.
+As evidenced by ratio () shown in Figure 3, the statistic generally assigns higher values to OOD samples than ID ones.
+However, the high variance of metric (Figure˜2) suggests the possibility of overoptimistic estimations.
 To address this issue, we introduce a correction term that exhibits an opposing behavior: it tends to be higher for ID samples than for OOD samples.
-Figure 5 shows that the perturbed activations of ID samples tend to be higher than those of OOD samples, especially in high-activation regions.
-We leverage this complementary signal by defining:
-where is a hyperparameter denoting the number of considered activations. We refine our OOD quantification by combining both metrics, weighted by a hyperparameter :
+Figure 5 in Appendix˜C shows that the perturbed activations of ID samples tend to be higher than those of OOD ones, especially in high-activation regions.
+We leverage this complementary signal by defining , where is a hyperparameter denoting the number of considered activations.
+We refine our OOD quantification by combining both metrics, weighted by a hyperparameter :
+Indeed, Figure˜3 illustrates that , suggesting that the correction term helps mitigate overconfident estimations.
 If be the set of values on ID validation samples, we could transform any into a normalized probability scale by constructing empirical cumulative distribution function (eCDF) derived from .
-The eCDF, denoted as , can be defined as:
-where is the indicator function.
+The eCDF, denoted as , can be defined as: , where is the indicator function.
 A higher value of indicates a higher likelihood of the sample being OOD.
-Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table 7).
-4.2 Revisiting Static Scaling Mechanisms
-Recent state-of-the-art post-processing methods for OOD detection, including SCALE [54], ASH [7], and LTS [8], operate by scaling either the activations or logits. The scaling factor is computed as:
-where denotes the percentile of all elements in activation .
-While this approach yields sample-specific scaling factors, it imposes a critical constraint: the percentile threshold is static and identical across all test samples, regardless of the nature of samples.
-We argue that this static nature limits the effectiveness of the scaling procedure and prevents optimal ID-OOD separability.
-Method ResNet-50 ResNet-101 RegNet-Y-16 ResNeXt-50 DenseNet-201 EfficientNetV2-L ViT-B-16 Swin-B Average near-OOD MSP 74.23 / 60.21 71.96 / 67.25 62.22 / 80.74 73.25 / 67.86 73.44 / 67.29 72.51 / 80.76 86.72 / 68.62 87.11 / 69.82 75.18 / 70.32 MLS 74.87 / 64.55 72.05 / 71.51 62.94 / 84.66 74.11 / 71.62 75.51 / 68.91 81.44 / 79.22 93.78 / 63.64 94.80 / 64.68 78.69 / 71.10 EBO 75.32 / 64.52 72.32 / 71.54 62.80 / 84.76 74.21 / 71.61 75.85 / 68.68 82.86 / 77.15 94.37 / 59.19 95.34 / 59.79 79.13 / 69.66 ReAct 72.61 / 68.81 68.07 / 75.00 70.73 / 75.37 70.96 / 74.13 69.97 / 73.65 72.36 / 71.39 86.63 / 68.35 82.64 / 73.26 74.25 / 72.50 ASH 69.47 / 71.33 65.24 / 76.61 82.51 / 67.81 70.98 / 75.25 92.83 / 52.30 94.85 / 44.78 94.45 / 53.20 96.37 / 47.58 83.34 / 61.11 SCALE 67.76 / 74.20 63.87 / 78.60 67.09 / 82.90 70.59 / 76.20 71.56 / 73.72 89.70 / 60.12 94.48 / 56.18 88.62 / 61.47 76.71 / 70.42 BFAct 72.35 / 68.88 67.96 / 75.16 78.72 / 66.09 70.96 / 74.14 71.20 / 72.61 75.53 / 62.46 82.09 / 70.66 71.81 / 75.28 73.83 / 70.66 LTS 68.01 / 73.37 63.91 / 78.27 69.82 / 80.75 70.27 / 76.20 71.29 / 74.56 87.30 / 73.63 88.83 / 67.43 86.61 / 67.22 75.76 / 73.93 OptFS 69.66 / 70.97 65.46 / 75.83 73.53 / 75.21 69.27 / 74.84 71.74 / 72.10 72.29 / 75.29 76.55 / 72.73 76.81 / 74.06 71.91 / 73.88 AdaSCALE-A 58.98 / 78.98 57.96 / 81.68 47.91 / 89.18 64.14 / 79.96 61.28 / 79.66 53.78 / 86.94 71.87 / 73.14 73.41 / 74.48 61.17 / 80.50 AdaSCALE-L 59.84 / 78.62 56.41 / 81.86 56.13 / 87.11 62.08 / 80.18 61.75 / 80.06 54.95 / 85.77 71.99 / 73.23 72.89 / 74.58 62.00 / 80.18 far-OOD MSP 53.15 / 84.06 53.87 / 83.81 40.41 / 90.08 53.07 / 84.21 53.60 / 84.43 54.74 / 87.92 56.41 / 84.62 73.39 / 82.02 54.83 / 85.14 MLS 42.57 / 88.19 43.89 / 88.30 32.92 / 93.70 44.91 / 87.97 48.43 / 87.44 68.64 / 84.80 81.89 / 81.42 95.16 / 73.37 57.30 / 85.65 EBO 42.72 / 88.09 44.30 / 88.23 32.47 / 93.82 45.12 / 87.86 48.95 / 87.15 74.48 / 81.13 86.95 / 76.34 96.08 / 63.99 58.88 / 83.33 ReAct 30.14 / 92.98 29.89 / 93.10 45.20 / 86.17 30.06 / 92.69 30.72 / 92.65 60.05 / 75.33 59.31 / 83.65 58.86 / 84.77 43.03 / 87.67 ASH 24.69 / 94.43 26.18 / 94.06 59.65 / 83.94 29.17 / 93.47 33.50 / 92.17 96.56 / 41.57 95.98 / 52.16 98.23 / 43.20 57.99 / 74.38 SCALE 21.44 / 95.39 22.54 / 95.05 32.16 / 94.16 30.62 / 93.54 33.17 / 92.70 89.63 / 62.58 88.36 / 72.32 86.59 / 66.77 50.56 / 84.06 BFAct 29.46 / 93.01 29.43 / 93.04 58.69 / 77.22 29.71 / 92.67 32.45 / 92.29 66.72 / 65.70 51.58 / 85.77 38.99 / 88.47 42.13 / 86.02 LTS 22.20 / 95.24 23.07 / 94.94 34.99 / 93.57 30.37 / 93.49 30.92 / 93.29 86.85 / 76.30 64.37 / 84.43 85.84 / 44.80 47.33 / 84.51 OptFS 25.66 / 93.87 26.97 / 93.55 47.37 / 86.73 27.54 / 93.40 34.42 / 91.04 53.62 / 83.62 46.11 / 87.35 44.27 / 87.79 38.25 / 89.67 AdaSCALE-A 17.84 / 96.14 18.51 / 95.95 21.37 / 95.84 22.08 / 95.24 28.01 / 93.23 37.61 / 91.48 47.63 / 86.83 47.81 / 87.14 30.11 / 92.73 AdaSCALE-L 17.92 / 96.12 19.15 / 95.76 20.10 / 96.19 22.16 / 95.01 28.00 / 93.18 38.81 / 90.51 47.28 / 86.97 46.24 / 87.97 29.96 / 92.71
+Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table˜9).
 4.3 Proposed Approach: Adaptive Scaling
 Building on our observations, we propose AdaSCALE (Adaptive SCALE), a novel approach that introduces dynamic, sample-specific adjustments to the scaling procedure.
-The key insight of AdaSCALE is that the percentile threshold should be a function of each test sample’s estimated OOD likelihood rather than a fixed value.
+The key insight is that percentile threshold should be a function of each test sample’s estimated OODness rather than a fixed value.
 The scaling factor increases as the percentile threshold rises (i.e., when more activations are excluded from the denominator in Equation 1).
 For optimal ID-OOD separation, we must scale ID samples more strongly than OOD samples, requiring a higher percentile for ID samples.
-Leveraging as an OOD likelihood estimator, we can now define an adaptive percentile threshold as:
-where and are hyperparameters that define the minimum and maximum limits of percentile threshold.
-It ensures that samples with lower OOD likelihood (i.e., likely ID samples) receive higher percentile thresholds, resulting in stronger scaling.
-Algorithm 1 summarizes our adaptive scaling factor computation procedure.
-We implement two variants: AdaSCALE-A scales activations as , similar to ASH [7] and SCALE [54].
-AdaSCALE-L scales logits as , similar to LTS [8].
-For OOD detection, we use the energy score on (directly or indirectly) scaled logits, with higher values indicating higher ID likelihood.
-This approach enables per-sample dynamic scaling, as outlined in Figure 2.
+We define an adaptive percentile threshold as: , where and are hyperparameters that define the minimum and maximum limits of percentile threshold.
+It ensures samples with lower estimated OODness receive higher percentile thresholds, resulting in stronger scaling.
+(See Algorithm˜1).
+We implement two variants: AdaSCALE-A scales activations as (Djurisic et al., 2023; Xu et al., 2024).
+AdaSCALE-L scales logits as (Djurisic et al., 2024).
+This approach is also outlined in Figure˜4.
 5 Experiments
@@
 For CIFAR experiments, we train three models per network using the standard cross-entropy loss and report the mean results across these three independent trials.
-The evaluation setup is provided in Table 2.
-Conventional OOD detection ID datasets Near-OOD Far-OOD Network CIFAR-10/100 CIFAR-100/10 MNIST, SVHN, WRN-28-10, DenseNet-101 TIN Textures, Places365 ImageNet-1k SSB-Hard iNaturalist, Places, EfficientNetV2-L, ResNet-101 NINCO OpenImage-O DenseNet-201, ViT-B-16 ImageNet-O Textures ResNet-50, ResNeXt-50 Places RegNet-Y-16, Swin-B Covariate shifted datasets for full spectrum OOD detection ImageNet-1k ImageNet-C, ImageNet-R, ImageNet-V2, ImageNet-ES
+The evaluation setup is provided in Table˜1.
+Although we tune for optimal results in each case (see Appendix˜G), near-optimal performance can be achieved by tuning only while fixing to across all 22 cases (See Table˜5).
+These values were determined via automatic parameter search (Yang et al., 2022; Zhang et al., 2023b) with respect to the AUROC metric using ResNet-50 model on ImageNet-1k benchmark.
+Best results are bold, and second-best results are underlined.
+Conventional OOD detection ID datasets Near-OOD Far-OOD Network CIFAR-10/100 CIFAR-100/10 MNIST, SVHN, WRN-28-10, DenseNet-101 TIN Textures, Places365 ImageNet-1k SSB-Hard iNaturalist, EfficientNetV2-L, ResNet-101 NINCO OpenImage-O DenseNet-201, ViT-B-16 ImageNet-O Textures ResNet-50, ResNeXt-50 Places RegNet-Y-16, Swin-B Covariate shifted datasets for full spectrum OOD detection ImageNet-1k ImageNet-C, ImageNet-R, ImageNet-V2, ImageNet-ES
 Metrics.
@@
 Baselines.
-We consider the following post-hoc methods: MSP [16], ODIN [30], EBO [33], ReAct [47], MLS [20], ASH [7], SCALE [54], BFAct [25], LTS [8], OptFS [64].
+MSP (Hendrycks & Gimpel, 2017), EBO (Liu et al., 2020), ReAct (Sun et al., 2021), MLS (Hendrycks et al., 2022a), ASH (Djurisic et al., 2023), SCALE (Xu et al., 2024), BFAct (Kong & Li, 2022), LTS (Djurisic et al., 2024), OptFS (Zhao et al., 2024).
 Currently, SCALE is the best-performing method (with ResNet-50), while OptFS is the best-generalizing method.
-Hyperparameters.
-The hyperparameters are determined via automatic parameter search provided by OpenOOD [57, 61].
-The best results are bold, and the second-best results are underlined across all results.
-We refer readers to Appendices A, B, C D, and E for complete details and studies.
 5.1 Empirical Results
-ImageNet-1k benchmark:
-The ImageNet-1k benchmark is crucial for OOD evaluation, as large-scale datasets offer a comprehensive representation of real-world scenarios.
-Consequently, post-hoc methods are particularly valuable when training data is inaccessible and retraining is computationally prohibitive.
-We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table 1.
+Method ResNet-50 ResNet-101 RegNet-Y-16 ResNeXt-50 DenseNet-201 EfficientNetV2-L ViT-B-16 Swin-B Average near-OOD MSP 74.23 / 60.21 71.96 / 67.25 62.22 / 80.74 73.25 / 67.86 73.44 / 67.29 72.51 / 80.76 86.72 / 68.62 87.11 / 69.82 75.18 / 70.32 MLS 74.87 / 64.55 72.05 / 71.51 62.94 / 84.66 74.11 / 71.62 75.51 / 68.91 81.44 / 79.22 93.78 / 63.64 94.80 / 64.68 78.69 / 71.10 EBO 75.32 / 64.52 72.32 / 71.54 62.80 / 84.76 74.21 / 71.61 75.85 / 68.68 82.86 / 77.15 94.37 / 59.19 95.34 / 59.79 79.13 / 69.66 ReAct 72.61 / 68.81 68.07 / 75.00 70.73 / 75.37 70.96 / 74.13 69.97 / 73.65 72.36 / 71.39 86.63 / 68.35 82.64 / 73.26 74.25 / 72.50 ASH 69.47 / 71.33 65.24 / 76.61 82.51 / 67.81 70.98 / 75.25 92.83 / 52.30 94.85 / 44.78 94.45 / 53.20 96.37 / 47.58 83.34 / 61.11 SCALE 67.76 / 74.20 63.87 / 78.60 67.09 / 82.90 70.59 / 76.20 71.56 / 73.72 89.70 / 60.12 94.48 / 56.18 88.62 / 61.47 76.71 / 70.42 BFAct 72.35 / 68.88 67.96 / 75.16 78.72 / 66.09 70.96 / 74.14 71.20 / 72.61 75.53 / 62.46 82.09 / 70.66 71.81 / 75.28 73.83 / 70.66 LTS 68.01 / 73.37 63.91 / 78.27 69.82 / 80.75 70.27 / 76.20 71.29 / 74.56 87.30 / 73.63 88.83 / 67.43 86.61 / 67.22 75.76 / 73.93 OptFS 69.66 / 70.97 65.46 / 75.83 73.53 / 75.21 69.27 / 74.84 71.74 / 72.10 72.29 / 75.29 76.55 / 72.73 76.81 / 74.06 71.91 / 73.88 AdaSCALE-A 58.98 / 78.98 57.96 / 81.68 47.91 / 89.18 64.14 / 79.96 61.28 / 79.66 53.78 / 86.94 71.87 / 73.14 73.41 / 74.48 61.17 / 80.50 AdaSCALE-L 59.84 / 78.62 56.41 / 81.86 56.13 / 87.11 62.08 / 80.18 61.75 / 80.06 54.95 / 85.77 71.99 / 73.23 72.89 / 74.58 62.00 / 80.18 far-OOD MSP 53.15 / 84.06 53.87 / 83.81 40.41 / 90.08 53.07 / 84.21 53.60 / 84.43 54.74 / 87.92 56.41 / 84.62 73.39 / 82.02 54.83 / 85.14 MLS 42.57 / 88.19 43.89 / 88.30 32.92 / 93.70 44.91 / 87.97 48.43 / 87.44 68.64 / 84.80 81.89 / 81.42 95.16 / 73.37 57.30 / 85.65 EBO 42.72 / 88.09 44.30 / 88.23 32.47 / 93.82 45.12 / 87.86 48.95 / 87.15 74.48 / 81.13 86.95 / 76.34 96.08 / 63.99 58.88 / 83.33 ReAct 30.14 / 92.98 29.89 / 93.10 45.20 / 86.17 30.06 / 92.69 30.72 / 92.65 60.05 / 75.33 59.31 / 83.65 58.86 / 84.77 43.03 / 87.67 ASH 24.69 / 94.43 26.18 / 94.06 59.65 / 83.94 29.17 / 93.47 33.50 / 92.17 96.56 / 41.57 95.98 / 52.16 98.23 / 43.20 57.99 / 74.38 SCALE 21.44 / 95.39 22.54 / 95.05 32.16 / 94.16 30.62 / 93.54 33.17 / 92.70 89.63 / 62.58 88.36 / 72.32 86.59 / 66.77 50.56 / 84.06 BFAct 29.46 / 93.01 29.43 / 93.04 58.69 / 77.22 29.71 / 92.67 32.45 / 92.29 66.72 / 65.70 51.58 / 85.77 38.99 / 88.47 42.13 / 86.02 LTS 22.20 / 95.24 23.07 / 94.94 34.99 / 93.57 30.37 / 93.49 30.92 / 93.29 86.85 / 76.30 64.37 / 84.43 85.84 / 44.80 47.33 / 84.51 OptFS 25.66 / 93.87 26.97 / 93.55 47.37 / 86.73 27.54 / 93.40 34.42 / 91.04 53.62 / 83.62 46.11 / 87.35 44.27 / 87.79 38.25 / 89.67 AdaSCALE-A 17.84 / 96.14 18.51 / 95.95 21.37 / 95.84 22.08 / 95.24 28.01 / 93.23 37.61 / 91.48 47.63 / 86.83 47.81 / 87.14 30.11 / 92.73 AdaSCALE-L 17.92 / 96.12 19.15 / 95.76 20.10 / 96.19 22.16 / 95.01 28.00 / 93.18 38.81 / 90.51 47.28 / 86.97 46.24 / 87.97 29.96 / 92.71
+Method CIFAR-10 CIFAR-100 Near-OOD Far-OOD Near-OOD Far-OOD MSP 43.18 / 89.07 34.49 / 90.88 55.64 / 80.23 61.73 / 76.82 MLS 51.54 / 89.33 39.62 / 91.68 57.24 / 81.25 60.19 / 78.92 EBO 51.54 / 89.37 39.58 / 91.75 57.45 / 81.10 60.12 / 78.96 ReAct 49.71 / 88.59 37.32 / 92.00 63.20 / 79.58 54.78 / 80.46 ASH 78.11 / 77.97 63.12 / 83.35 80.97 / 70.09 69.38 / 79.06 SCALE 53.00 / 89.20 39.27 / 91.93 58.38 / 81.00 57.19 / 80.56 BFAct 54.90 / 88.56 43.05 / 90.66 72.26 / 74.70 57.44 / 77.63 LTS 55.71 / 88.77 41.06 / 91.74 59.98 / 80.60 80.48 / 81.79 OptFS 64.82 / 85.72 47.67 / 89.99 76.80 / 73.02 60.23 / 77.76 AdaSCALE-A 43.07 / 90.31 33.11 / 92.66 57.33 / 81.35 54.53 / 81.14 AdaSCALE-L 44.71 / 90.14 33.43 / 92.69 58.70 / 81.07 52.49 / 82.21
+CIFAR benchmark.
+We compare AdaSCALE with post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance in Table˜3.
+AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection.
+For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points.
+Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive.
+In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21.
+While activation-shaping methods perform well in ImageNet-1k, they seem to underperform in CIFAR.
+In contrast, AdaSCALE achieves consistently superior performance across all setups.
+ImageNet-1k benchmark.
+We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table˜2.
 AdaSCALE demonstrates consistently strong performance across all architectures compared to existing methods.
@@
 In contrast, AdaSCALE-A achieves the best performance in near-OOD detection across all architectures, except for Swin-B, where BFAct performs optimally.
-Furthermore, the effectiveness of AdaSCALE extends beyond the near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% in the FPR@95 metric compared to OptFS.
-Method Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC MSP 86.75 50.13 74.28 65.67 MLS 88.93 49.48 76.33 65.06 EBO 89.11 49.08 77.21 63.44 ReAct 87.22 51.38 67.23 69.53 ASH 87.01 52.02 72.36 65.70 SCALE 86.75 52.27 69.36 68.97 BFAct 87.12 51.14 66.13 69.69 LTS 86.46 53.29 66.76 71.63 OptFS 85.83 52.17 63.32 71.44 AdaSCALE-A 81.34 55.03 58.87 72.41 AdaSCALE-L 81.62 55.14 59.19 72.85
-Full-Spectrum OOD Detection:
-Full-spectrum OOD (FS-OOD) detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs.
-We present FS-OOD detection results in Table 3.
-We can observe that this is a highly challenging task, as covariate-shifted ID datasets cause a significant performance drop for all methods compared to the conventional case.
-Despite this, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in the FPR@95 metric for full-spectrum OOD detection across both near- and far-OOD datasets.
-ISH training:
-Apart from enhancing the prior postprocessor ASH [7], SCALE [56] introduces a training regularization to emphasize samples with more distinct ID characteristics.
-We assess the performance of each method in ResNet-50 and ResNet-101 model following this regularization in Table 4.
-The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by and in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively.
-Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in a full-spectrum setting.
-Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%.
-These findings also seem to generalize to the ResNet-101 network, as shown in Table 4.
-Method OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD ResNet-50 MSP 74.07 / 62.16 51.13 / 84.64 87.52 / 40.36 74.41 / 61.52 MLS 74.38 / 66.43 41.57 / 88.90 88.89 / 39.49 71.53 / 61.69 EBO 74.68 / 66.46 41.85 / 88.83 89.05 / 39.18 71.77 / 61.11 ReAct 71.98 / 70.81 28.76 / 93.49 87.78 / 43.88 61.87 / 71.64 ASH 67.99 / 73.46 23.88 / 94.67 85.74 / 45.29 57.81 / 72.81 SCALE 65.68 / 76.41 20.77 / 95.62 84.31 / 48.40 54.48 / 74.79 BFAct 71.59 / 70.85 28.38 / 93.50 87.51 / 43.95 61.39 / 71.43 LTS 66.32 / 75.03 22.07 / 95.28 85.08 / 46.16 57.11 / 73.06 OptFS 67.71 / 73.03 24.65 / 94.18 85.38 / 45.91 57.09 / 72.95 AdaSCALE-A 57.43 / 80.86 16.52 / 96.46 80.85 / 51.24 51.57 / 75.63 AdaSCALE-L 56.83 / 80.81 17.62 / 96.22 80.97 / 50.59 53.43 / 74.62 ResNet-101 MSP 71.39 / 68.31 51.00 / 84.81 85.70 / 45.72 73.69 / 62.79 MLS 72.32 / 71.94 41.04 / 88.99 87.39 / 44.88 69.94 / 63.23 EBO 72.78 / 71.92 41.45 / 88.87 87.64 / 44.66 70.23 / 62.65 ReAct 67.74 / 75.74 28.53 / 93.47 85.40 / 48.83 60.50 / 72.18 ASH 66.03 / 77.79 25.21 / 94.43 83.95 / 50.57 56.91 / 73.37 SCALE 64.30 / 78.98 23.09 / 94.95 83.14 / 51.05 55.88 / 73.27 BFAct 67.53 / 75.86 28.32 / 93.40 85.11 / 48.96 60.12 / 71.83 LTS 66.32 / 75.03 22.07 / 95.28 85.08 / 46.16 57.11 / 73.06 OptFS 67.71 / 73.03 24.65 / 94.18 85.38 / 45.91 57.09 / 72.95 AdaSCALE-A 54.66 / 83.52 16.81 / 96.32 78.52 / 55.19 49.92 / 76.20 AdaSCALE-L 53.91 / 83.49 17.55 / 96.15 78.63 / 54.60 51.47 / 75.47
-CIFAR benchmark:
-CIFAR benchmarks are traditionally used to evaluate OOD detectors in smaller-scale settings.
-We compare AdaSCALE with recent state-of-the-art post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance over both architectures in Table 5.
-AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection.
-For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points.
-Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive.
-In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21.
-While activation-shaping methods (ReAct, ASH, SCALE, OptFS) perform remarkably well in comparison to logit-based methods in a large-scale setting (ImageNet-1k), they generally seem to underperform in small-scale settings (CIFAR).
-In contrast, AdaSCALE’s achieves consistently superior performance across all setups, demonstrating its robustness.
-Method CIFAR-10 CIFAR-100 Near-OOD Far-OOD Near-OOD Far-OOD MSP 43.18 / 89.07 34.49 / 90.88 55.64 / 80.23 61.73 / 76.82 MLS 51.54 / 89.33 39.62 / 91.68 57.24 / 81.25 60.19 / 78.92 EBO 51.54 / 89.37 39.58 / 91.75 57.45 / 81.10 60.12 / 78.96 ReAct 49.71 / 88.59 37.32 / 92.00 63.20 / 79.58 54.78 / 80.46 ASH 78.11 / 77.97 63.12 / 83.35 80.97 / 70.09 69.38 / 79.06 SCALE 53.00 / 89.20 39.27 / 91.93 58.38 / 81.00 57.19 / 80.56 BFAct 54.90 / 88.56 43.05 / 90.66 72.26 / 74.70 57.44 / 77.63 LTS 55.71 / 88.77 41.06 / 91.74 59.98 / 80.60 80.48 / 81.79 OptFS 64.82 / 85.72 47.67 / 89.99 76.80 / 73.02 60.23 / 77.76 AdaSCALE-A 43.07 / 90.31 33.11 / 92.66 57.33 / 81.35 54.53 / 81.14 AdaSCALE-L 44.71 / 90.14 33.43 / 92.69 58.70 / 81.07 52.49 / 82.21
-Accuracy:
-5.2 Ablation / hyperparameter studies
-Sensitivity of , , and .
-As discussed in Sec.
-4.1.2, ID and OOD distinctions are more pronounced in high activations but diminish as more activations are included.
-We analyze the impact of (used in activation shift) and (used in perturbed activation) using a heatmap of validation AUROC with ResNet-50 (Figure 6(a)).
-The darker region indicates higher AUROC, suggesting optimal values of and for ResNet-50 model.
-Furthermore, the heatmap suggests that is far more critical hyperparameter than .
-The hyperparameter controls the weighting of in computing , the predetermined OOD likelihood.
-The sensitivity analysis is presented in Figure 6(b) which shows near-OOD and far-OOD detection using FPR@95.
-It suggests optimal FPR@95 is achieved at .
-Predetermined OOD likelihood .
-Adaptive scaling depends on predetermined OOD likelihood to determine the extent of scaling.
-We study the effect of various predetermined OOD likelihood functions on OOD detection using ResNet-50 network (ImageNet-1k) in Table 6.
+Furthermore, effectiveness of AdaSCALE extends beyond near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% over OptFS in the FPR@95 metric.
+Method Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC ReAct 87.22 51.38 67.23 69.53 ASH 87.01 52.02 72.36 65.70 SCALE 86.75 52.27 69.36 68.97 BFAct 87.12 51.14 66.13 69.69 LTS 86.46 53.29 66.76 71.63 OptFS 85.83 52.17 63.32 71.44 AdaSCALE-A 81.34 55.03 58.87 72.41 AdaSCALE-L 81.62 55.14 59.19 72.85
+FSOOD Detection.
+FSOOD detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs.
+We present FSOOD detection results in Table˜4.
+We can observe that this is a highly challenging task, as covariate-shifted ID datasets cause a significant performance drop for all methods compared to conventional case.
+Despite this, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in the FPR@95 metric for FSOOD detection across both near- and far-OOD datasets.
+Accuracy: Like SCALE and LTS, AdaSCALE applies linear transformations to scale activations or logits, preserving accuracy, unlike post-hoc rectification methods (ASH, ReAct).
+5.2 Ablation / hyperparameter studies / Discussion
+Sensitivity study.
+We study the sensitivity of AdaSCALE-A hyperparameters with ResNet-50 on ImageNet-1k (Table˜7).
+The optimal perturbation magnitude is , close to RGB means and sufficient to disturb peak activations.
+Setting ensures activation shift dominates the regularizer .
+Optimal performance at confirms that shifts in peak activations are the most reliable OOD signal, while empirically performs best.
+yielding optimal result supports our hypothesis that perturbation should be trivial enough to only disturb peak activations.
+(See Appendix˜E for complete sensitivity study.)
+Method ImageNet-1k CIFAR-100 Near-OOD Far-OOD Near-OOD Far-OOD OptFS 71.91 / 73.88 38.25 / 89.67 76.81 / 73.02 60.23 / 77.76 AdaSCALE-A 62.29 / 79.72 32.72 / 91.82 58.00 / 81.14 56.47 / 80.99
+Constrained hyperparameter tuning.
+The hyperparameters = determined with ResNet-50 model in ImageNet-1k benchmark generalize across architectures and datasets.
+To demonstrate this, we conduct experiments only allowing to be tuned across 8 and 2 diverse architectures in ImageNet-1k and CIFAR-100 datasets, respectively and present the results in Table˜5.
+Inspired by early works that tune the percentile hyperparameter in , we set to the lower limit 60.
+Even under this highly constrained setting, AdaSCALE significantly outperforms the previous state-of-the-art (OptFS) by an average of 13% (FPR95) / 8% (AUROC) on near-OOD and 14% (FPR95) / 2% (AUROC) on far-OOD datasets on ImageNet-1k benchmark.
+The superiority of AdaSCALE over OptFS is also evident in a small-scale setting (CIFAR-100).
+Estimated OODness .
+Adaptive scaling depends on estimated OODness to determine the extent of scaling.
+We study the effect of various estimated OODness functions using ResNet-50 network (ImageNet-1k) in Table˜6.
 It clearly shows Q component of being most critical while as correction term being a relatively superior choice.
-OOD Detection Near-OOD Far-OOD Q 59.43 / 78.14 19.70 / 95.73 70.39 / 74.00 21.40 / 95.31 58.97 / 78.98 17.84 / 96.14 65.11 / 76.23 19.76 / 95.67 58.91 / 78.74 18.02 / 96.08
-Adaptive percentile.
-Unlike SCALE which uses constant percentile, AdaSCALE uses dynamic percentile lying in range adaptive to each sample.
-We show the effect of various percentile limit ranges in Figure 7 in the form of a heatmap on the AUROC validation metric on ImageNet-1k benchmark.
-The extent of darkness in the heatmap conveys a strong performance (corresponding to highest AUROC / lowest FPR@95).
-The diagonal entries, representing the results of SCALE, are lighter in comparison to the rest of the cells, denoting the results of AdaSCALE.
-Hence, it can be observed that using adaptive percentile leads to better OOD detection in comparison to static percentile.
+However, estimated OODness alone – without adaptive scaling – does not result in strong performance.
+Near-OOD Far-OOD without scaling 79.81 / 72.32 84.00 / 68.13 59.43 / 78.14 19.70 / 95.73 70.39 / 74.00 21.40 / 95.31 58.97 / 78.98 17.84 / 96.14 65.11 / 76.23 19.76 / 95.67 58.91 / 78.74 18.02 / 96.08
+| Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | |||||
+| 0.1 | 63.76 / 77.50 | 19.26 / 95.85 | 1 | 67.49 / 75.45 | 20.41 / 95.54 | 1% | 58.97 / 78.98 | 17.84 / 96.14 | 1% | 59.08 / 78.56 | 18.35 / 96.02 | 1% | 61.77 / 78.29 | 19.28 / 95.77 |
+| 0.5 | 58.97 / 78.98 | 17.84 / 96.14 | 10 | 58.97 / 78.98 | 17.84 / 96.14 | 10% | 60.40 / 77.89 | 19.72 / 95.63 | 5% | 58.97 / 78.98 | 17.84 / 96.14 | 5% | 58.97 / 78.98 | 17.84 / 96.14 |
+| 1.0 | 61.60 / 76.96 | 19.31 / 95.84 | 100 | 59.03 / 78.10 | 19.39 / 95.80 | 100% | 60.99 / 76.49 | 21.88 / 94.96 | 100% | 62.43 / 75.44 | 23.05 / 94.69 | 100% | 67.37 / 73.08 | 22.93 / 95.01 |
+| Method | Gaussian blur | JPEG compression | Adversarial training | |||
+| Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | |
+| SCALE | 72.68 / 70.15 | 29.65 / 92.72 | 69.73 / 71.84 | 27.46 / 93.56 | 66.99 / 76.10 | 24.77 / 94.40 |
+| OptFS | 71.84 / 69.66 | 30.27 / 92.24 | 69.43 / 71.12 | 28.34 / 92.95 | 70.11 / 73.51 | 28.32 / 93.42 |
+| AdaSCALE-A | 67.53 / 75.18 | 28.11 / 93.31 | 61.24 / 77.59 | 25.87 / 94.06 | 60.21 / 80.68 | 21.13 / 95.35 |
+Robustness to real-world image corruptions.
+To evaluate the practical robustness of our method, we assess its performance when input images are subjected to common real-world corruptions.
+We compare AdaSCALE against SCALE and OptFS on the ImageNet-1k dataset using a ResNet-50, applying Gaussian blur (kernel=5, sigma=(0.01, 0.5)) and JPEG compression (50%) to the input images.
+The results (FPR@95/AUROC), presented in Table˜8 without any hyperparameter re-tuning, show that AdaSCALE’s performance advantage is maintained even under these conditions.
+Robustness to adversarial training.
+We evaluate AdaSCALE’s performance on adversarially trained model (=0.05) using the exact same hyperparameters as for the standard ResNet-50.
+The results (FPR@95 / AUROC) in Table˜8 show that AdaSCALE outperforms baseline methods (SCALE and OptFS).
+This highlights the robustness of AdaSCALE across diverse training schemes.
+Near-OOD Far-OOD 10 59.69 / 78.52 18.25 / 96.03 100 59.05 / 78.92 17.79 / 96.13 1000 58.99 / 78.95 17.86 / 96.13 5000 58.97 / 78.98 17.84 / 96.14
 ID statistics.
 With the rise of large models, where training data is often undisclosed or inaccessible, relying on full training ID datasets for OOD detection has become increasingly impractical.
-While our approach leverages ID statistics for optimal performance, we rigorously assess its effectiveness with limited data.
-We conduct experiments on ImageNet-1k using ID samples to compute ID statistics, where .
-The results in Table 7 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
-OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD 10 59.69 / 78.52 18.25 / 96.03 81.66 / 49.76 52.18 / 75.01 100 59.05 / 78.92 17.79 / 96.13 81.52 / 49.19 52.39 / 74.82 1000 58.99 / 78.95 17.86 / 96.13 81.55 / 49.48 52.24 / 74.94 5000 58.97 / 78.98 17.84 / 96.14 81.52 / 49.35 52.33 / 74.89
+We rigorously assess AdaSCALE’s effectiveness with limited data by conducting experiments on ImageNet-1k with ResNet-50 using ID samples to compute ID statistics, where .
+The results in Table˜9 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
+Fixed percentile (SCALE/LTS) 33 40 45 48 54 Variable percentile (AdaSCALE) 152 149 155 152 164 Latency ratio (AdaSCALE / SCALE) 4.66 3.76 3.42 3.14 3.02
+Latency.
+AdaSCALE incurs extra forward pass to compute perturbed activation .
+Also, top-k operations (time complexity: ) are applied to and to estimate OODness.
+Comparing variable vs.
+fixed percentiles for scaling in Table˜10 over 10,000 trials, we observe that variable percentiles induce higher latency, though the latency ratio decreases with higher-dimensional activation spaces.
+While SCALE and OptFS have similar latency, AdaSCALE’s is 2.91x higher with gradient attribution for trivial pixel perturbation, dropping to 1.56x with random pixel perturbation.
 6 Conclusion
-We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts the scaling process based on a sample’s estimated OOD likelihood.
+We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts the scaling process based on a sample’s estimated OODness.
 Leveraging the observation that OOD samples exhibit larger activation shifts under minor perturbations, AdaSCALE assigns stronger scaling to likely ID samples and weaker scaling to likely OOD samples, enhancing ID-OOD separability.
 AdaSCALE achieves state-of-the-art performance as well as generalization across architectures requiring negligibly few ID samples, making it highly practical for real-world deployment.
-Supplementary Material
+It lays the foundation for future research to build upon the adaptive scaling mechanism.
+7 Broader Impacts
+This work has positive impact in trustworthy deep learning by enabling detection of OOD samples.
+Reliable OOD detection can help prevent models from making overconfident predictions in high-stakes applications.
+Appendix
 Appendix A Notations
@@
 Rectified Linear Unit activation function: .
-Appendix B Additional studies
-Image Perturbation.
-A sufficiently small perturbation, as discussed in Sec.
-4.3, is used to make distinctive observations in ID vs OOD.
-We now systematically investigate the impact of the extent and nature of perturbation on pre-determining OOD likelihood which is in-turn responsible for OOD detection.
-We present the results in Table 9.
-It is clearly evident that perturbing regions with trivial intensity value leads to better OOD detection.
-We can infer that with such perturbation, acts as a more reliable proxy for OOD likelihood.
-Another key takeaway is that randomly perturbing small regions can also lead to strong performance.
-This finding highlights a more efficient approach to OOD detection, relying solely on minor random perturbations rather than requiring a predetermination of trivial pixels.
-Furthermore, the sensitivity study of presented at Table 10 suggests the optimal value of to be around 0.5.
-Region type OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD Random 61.73 / 78.15 19.44 / 95.74 83.19 / 48.59 53.45 / 74.10 59.97 / 78.67 18.14 / 96.06 81.92 / 49.19 52.35 / 74.84 60.27 / 78.02 18.45 / 96.00 82.07 / 48.62 52.84 / 74.70 62.81 / 76.27 19.95 / 95.70 83.40 / 46.55 54.94 / 73.42 Trivial 61.77 / 78.29 19.28 / 95.77 82.92 / 48.86 53.34 / 74.24 58.97 / 78.98 17.84 / 96.14 81.52 / 49.35 52.33 / 74.89 60.24 / 78.17 17.94 / 96.08 82.19 / 48.59 52.59 / 74.77 66.43 / 74.10 21.58 / 95.29 85.26 / 44.86 56.56 / 72.82 Significant 69.43 / 75.13 22.62 / 95.17 85.59 / 48.32 53.76 / 75.23 67.31 / 75.78 21.24 / 95.44 85.07 / 48.40 53.26 / 75.63 65.65 / 76.20 20.39 / 95.62 84.36 / 48.33 53.17 / 75.55 64.54 / 75.39 20.48 / 95.61 83.78 / 47.07 54.11 / 74.54 All 67.37 / 73.08 22.93 / 95.01 85.66 / 44.00 57.80 / 72.19
-Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC 0.1 63.76 77.50 19.26 95.85 0.5 58.97 78.98 17.84 96.14 1.0 61.60 76.96 19.31 95.84
-Latency.
-AdaSCALE incurs extra computational overhead in comparison to its preceding works [7, 54, 8].
-It uses an extra forward pass to obtain perturbed activation .
-On top of that, top-k operations (time complexity: , where D is dimension of activation space) on and are employed for pre-determining OOD likelihood.
-Furthermore, comparing variable percentile to fixed percentile for scaling in Table 11 across 10000 trials, we observe variable percentile inducing relatively higher latency.
-The latency ratio decreases as activation space with a higher dimension is considered.
-All experiments are performed using NVIDIA RTX 4090 GPUs.
-Fixed percentile (SCALE) 33 40 45 48 54 Variable percentile (AdaSCALE) 152 149 155 152 164 Latency ratio (AdaSCALE / SCALE) 4.66 3.76 3.42 3.14 3.02
-Appendix C Hyperparameters
-All hyperparameters are determined with respect to the AUROC metric using automatic parameter search of OpenOOD [57, 61].
-Although AdaSCALE may appear to require numerous hyperparameters, our findings indicate that setting to consistently yields near-optimal performance across all setups.
-Consequently, it can be inferred that only the hyperparameters and need to be appropriately tuned for any new architecture for near-optimal performance.
-| Dataset | Network | Hyperparameters | ||||||
-| CIFAR-10 | WideResNet-28-10 | 60 | 95 | 10 | 0.5 | |||
-| DenseNet-101 | 65 | 90 | 10 | 0.5 | ||||
-| CIFAR-100 | WideResNet-28-10 | 60 | 85 | 10 | 0.5 | |||
-| DenseNet-101 | 70 | 80 | 10 | 0.5 | ||||
-| ImageNet-1k | ResNet-50 | 80 | 85 | 10 | 0.5 | |||
-| ResNet-101 | 80 | 85 | 10 | 0.5 | ||||
-| RegNet-Y-16 | 60 | 90 | 10 | 0.5 | ||||
-| ResNeXt-50 | 80 | 85 | 10 | 0.5 | ||||
-| DenseNet-201 | 90 | 95 | 10 | 0.5 | ||||
-| EfficientNetV2-L | 60 | 99 | 10 | 0.5 | ||||
-| Vit-B-16 | 60 | 85 | 10 | 0.5 | ||||
-| Swin-B | 90 | 99 | 10 | 0.5 | ||||
-| Dataset | Network | Hyperparameters | ||||||
-| CIFAR-10 | WideResNet-28-10 | 60 | 85 | 10 | 0.5 | |||
-| DenseNet-101 | 70 | 85 | 10 | 0.5 | ||||
-| CIFAR-100 | WideResNet-28-10 | 60 | 80 | 10 | 0.5 | |||
-| DenseNet-101 | 65 | 75 | 10 | 0.5 | ||||
-| ImageNet-1k | ResNet-50 | 80 | 85 | 10 | 0.5 | |||
-| ResNet-101 | 70 | 80 | 10 | 0.5 | ||||
-| RegNet-Y-16 | 60 | 85 | 10 | 0.5 | ||||
-| ResNeXt-50 | 70 | 80 | 10 | 0.5 | ||||
-| DenseNet-201 | 90 | 95 | 10 | 0.5 | ||||
-| EfficientNetV2-L | 60 | 99 | 10 | 0.5 | ||||
-| Vit-B-16 | 75 | 85 | 10 | 0.5 | ||||
-| Swin-B | 90 | 99 | 10 | 0.5 | ||||
-Appendix D CIFAR-results
-D.1 WRN-28-10
+Appendix B Algorithm
+The algorithm for computing adaptive scaling factor is provided in Algorithm˜1.
+Appendix C Additional observation in activation space.
+Figure˜5 shows perturbed activations are, on average, higher for ID samples than for OOD samples.
+Appendix D Additional Experiments and Analysis
+D.1 Superiority of the Proposed OODness Metric ()
+To validate our choice of the OODness metric, we conducted an extensive ablation study comparing the performance of our proposed metric, against alone.
+The results, summarized in Table 12, demonstrate that using provides a consistent and significant performance improvement over using alone.
+This advantage holds across all eight diverse architectures.
+| Category | Metric | ResNet-50 | ResNet-101 | RegNet-Y-16 | ResNeXt-50 | DenseNet-201 | EfficientNetV2-L | ViT-B-16 | Swin-B | Average |
+| near-OOD | 59.65/77.82 | 56.67/81.65 | 60.63/86.00 | 63.99/79.34 | 68.65/78.03 | 73.18/80.17 | 82.61/69.31 | 75.85/72.26 | 68.90/78.07 | |
+| 58.98/78.98 | 57.96/81.68 | 47.91/89.18 | 64.14/79.96 | 61.28/79.66 | 53.78/86.94 | 71.87/73.14 | 73.41/74.48 | 61.17/80.50 | ||
+| far-OOD | 19.82/95.72 | 20.32/95.50 | 31.53/93.94 | 26.87/94.32 | 35.55/92.19 | 66.20/84.60 | 77.52/76.39 | 63.20/80.78 | 42.63/89.18 | |
+| 17.84/96.14 | 18.51/95.95 | 21.37/95.84 | 22.08/95.24 | 28.01/93.23 | 37.61/91.48 | 47.63/86.83 | 47.81/87.14 | 30.11/92.73 |
+For instance, if an OOD sample gets a very high score, it will produce a very large scaling factor.
+This can amplify the logits so much that it washes out the subtle but important information contained within the values of the logits themselves.
+By adding the opposing term, we "tame" these potentially overconfident scaling factors.
+This ensures a healthier balance where both the feature-level instability (captured by ) and the original logit distribution contribute to the final OOD score.
+Illustrative Example: A simple toy example demonstrates why this balancing is critical.
+Logits: Let and .
+Case 1: Overconfident Scaling (using only )
+Let’s say this yields scaling factors and . Assume a threshold of 25, where energy > 25 is ID and energy < 25 is OOD (We deal with magnitude for simplicity).
+Here, , leading to an OOD detection failure.
+Case 2: Balanced Scaling (using to temper the scaling)
+This leads to less extreme scaling, e.g., and . Assume a threshold of 5, where energy > 5.5 is ID and energy < 5.5 is OOD (We deal with magnitude for simplicity).
+Here, , leading to a successful OOD detection.
+D.2 Comparison with Distance-Based OOD Detection Methods
+To situate AdaSCALE within the broader landscape of OOD detection techniques, we compare its performance against representative distance-based methods, specifically NNGuide and MDS.
+The experiments were conducted on the large-scale ImageNet-1k benchmark using a ResNet-50 backbone.
+As shown in Table 13, AdaSCALE demonstrates a significant performance advantage over these methods, particularly in the challenging near-OOD setting.
+| Method | Near-OOD | Far-OOD | ||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| NNGuide | 72.46 | 68.09 | 21.40 | 95.33 |
+| MDS | 76.27 | 64.53 | 70.82 | 66.22 |
+| AdaSCALE-A | 58.97 | 78.98 | 17.84 | 96.14 |
+D.3 Analysis on Adversarially Trained Models
+We investigate the effectiveness of our approach on adversarially trained models to test its robustness.
+This analysis is twofold: first, we verify that the core assumption of our method holds, and second, we evaluate its OOD detection performance.
+D.3.1 Activation Sensitivity in Adversarially Trained Networks
+The effectiveness of our method relies on the activation shift ratio () being greater than 1 for OOD samples compared to ID samples.
+We measured this ratio by taking the top-1% activation shifts on a ResNet-50 model adversarially trained on ImageNet-1k.
+The results in Table 14 confirm that the ratio remains consistently greater than 1 across various OOD datasets.
+This demonstrates that OOD samples still produce a larger activation shift than ID samples, even in adversarially robust models, validating our method’s applicability.
+| OOD Dataset | () | () |
+| SSB-hard | 1.78 | 1.73 |
+| NINCO | 1.29 | 1.28 |
+| ImageNet-O | 1.69 | 1.72 |
+| OpenImage-O | 1.48 | 1.45 |
+| iNaturalist | 1.14 | 1.14 |
+| Textures | 1.33 | 1.18 |
+| Places | 1.30 | 1.16 |
+D.3.2 Performance of AdaSCALE on Adversarially Trained Networks
+We evaluate AdaSCALE’s OOD detection performance on these adversarially trained models using the exact same hyperparameters as for the standard ResNet-50.
+The results in Table 15 show that AdaSCALE consistently outperforms baseline methods.
+This highlights the robustness of our approach AdaSCALE, which remains superior even when applied to models with different training schemes without any specific hyperparameter tuning.
+| Training | Method | Near-OOD | Far-OOD | ||
+| FPR@95 | AUROC | FPR@95 | AUROC | ||
+| SCALE | 68.54 | 75.43 | 25.95 | 94.40 | |
+| OptFS | 70.36 | 72.97 | 28.60 | 93.24 | |
+| AdaSCALE-A | 60.73 | 80.84 | 19.90 | 95.74 | |
+| SCALE | 66.99 | 76.10 | 24.77 | 94.40 | |
+| OptFS | 70.11 | 73.51 | 28.32 | 93.42 | |
+| AdaSCALE-A | 60.21 | 80.68 | 21.13 | 95.35 | |
+D.4 Robustness to Real-World Image Corruptions
+To evaluate the practical robustness of our method, we assess its performance when input images are subjected to common real-world corruptions.
+We benchmark AdaSCALE against SCALE and OptFS on the ImageNet-1k dataset using a ResNet-50, applying Gaussian blur and JPEG compression to the input images.
+The results, presented without any hyperparameter re-tuning, show that AdaSCALE’s performance advantage is maintained even under these challenging conditions.
+| Method | Near-OOD | Far-OOD | ||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| SCALE | 72.68 | 70.15 | 29.65 | 92.72 |
+| OptFS | 71.84 | 69.66 | 30.27 | 92.24 |
+| AdaSCALE-A | 67.53 | 75.18 | 28.11 | 93.31 |
+| Method | Near-OOD | Far-OOD | ||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| SCALE | 69.73 | 71.84 | 27.46 | 93.56 |
+| OptFS | 69.43 | 71.12 | 28.34 | 92.95 |
+| AdaSCALE-A | 61.24 | 77.59 | 25.87 | 94.06 |
+Appendix E Sensitivity study
+E.1 Sensitivity study of
+The sensitivity study of presented at Table˜18 suggests the optimal value of to be around 0.5.
+| Near-OOD | Far-OOD | |||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| 0.1 | 63.76 | 77.50 | 19.26 | 95.85 |
+| 0.5 | 58.97 | 78.98 | 17.84 | 96.14 |
+| 1.0 | 61.60 | 76.96 | 19.31 | 95.84 |
+E.2 Sensitivity study of
+The sensitivity study of presented at Table˜18 suggests the optimal value of to be around 10.
+| Near-OOD | Far-OOD | |||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| 0.1 | 70.07 | 74.12 | 21.40 | 95.33 |
+| 1 | 67.49 | 75.45 | 20.41 | 95.54 |
+| 10 | 58.97 | 78.98 | 17.84 | 96.14 |
+| 100 | 59.03 | 78.10 | 19.39 | 95.80 |
+E.3 Sensitivity study of
+The sensitivity study of presented at Table˜20 suggests the optimal value of to be around 1%.
+| Near-OOD | Far-OOD | |||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| 1% | 58.97 | 78.98 | 17.84 | 96.14 |
+| 5% | 60.36 | 77.74 | 19.95 | 95.63 |
+| 10% | 60.40 | 77.89 | 19.72 | 95.63 |
+| 50% | 60.34 | 77.25 | 20.90 | 95.25 |
+| 100% | 60.99 | 76.49 | 21.88 | 94.96 |
+E.4 Sensitivity study of
+The sensitivity study of presented at Table˜21 suggests the optimal value of to be around 5%.
+| Near-OOD | Far-OOD | |||
+| FPR@95 | AUROC | FPR@95 | AUROC | |
+| 1% | 59.08 | 78.56 | 18.35 | 96.02 |
+| 5% | 58.97 | 78.98 | 17.84 | 96.14 |
+| 10% | 59.41 | 78.65 | 18.33 | 95.96 |
+| 100% | 62.43 | 75.44 | 23.05 | 94.69 |
+E.5 Sensitivity study of
+We present the complete results of image perturbation study (FPR@95 / AUROC ) in Table˜22.
+| Pixel type | OOD Detection | FS-OOD Detection | |||
+| Near-OOD | Far-OOD | Near-OOD | Far-OOD | ||
+| Random | 61.73 / 78.15 | 19.44 / 95.74 | 83.19 / 48.59 | 53.45 / 74.10 | |
+| 59.97 / 78.67 | 18.14 / 96.06 | 81.92 / 49.19 | 52.35 / 74.84 | ||
+| 60.27 / 78.02 | 18.45 / 96.00 | 82.07 / 48.62 | 52.84 / 74.70 | ||
+| 62.81 / 76.27 | 19.95 / 95.70 | 83.40 / 46.55 | 54.94 / 73.42 | ||
+| Trivial | 61.77 / 78.29 | 19.28 / 95.77 | 82.92 / 48.86 | 53.34 / 74.24 | |
+| 58.97 / 78.98 | 17.84 / 96.14 | 81.52 / 49.35 | 52.33 / 74.89 | ||
+| 60.24 / 78.17 | 17.94 / 96.08 | 82.19 / 48.59 | 52.59 / 74.77 | ||
+| 66.43 / 74.10 | 21.58 / 95.29 | 85.26 / 44.86 | 56.56 / 72.82 | ||
+| Salient | 69.43 / 75.13 | 22.62 / 95.17 | 85.59 / 48.32 | 53.76 / 75.23 | |
+| 67.31 / 75.78 | 21.24 / 95.44 | 85.07 / 48.40 | 53.26 / 75.63 | ||
+| 65.65 / 76.20 | 20.39 / 95.62 | 84.36 / 48.33 | 53.17 / 75.55 | ||
+| 64.54 / 75.39 | 20.48 / 95.61 | 83.78 / 47.07 | 54.11 / 74.54 | ||
+| All | 67.37 / 73.08 | 22.93 / 95.01 | 85.66 / 44.00 | 57.80 / 72.19 | |
+E.6 Sensitivity study of
+We present sensitivity study (FPR@95 / AUROC ) of , setting to 85 in the Table˜23.
+| 70 | 75 | 80 | 85 | |
+| Near-OOD | 62.54 / 79.60 | 58.57 / 81.68 | 58.98 / 78.98 | 67.76 / 74.20 |
+| Far-OOD | 21.62 / 95.19 | 19.86 / 95.62 | 17.84 / 96.14 | 21.44 / 95.39 |
+E.7 Joint sensitivity study
+We present joint sensitivity study in terms of near-OOD detection among three hyperparameters (, below:
+| \ | 0.1 | 1 | 10 | 100 |
+| 1% | 70.07 / 74.12 | 67.49 / 75.45 | 58.97 / 78.98 | 59.03 / 78.10 |
+| 5% | 69.53 / 74.33 | 63.87 / 76.84 | 60.36 / 77.74 | 61.12 / 77.22 |
+| 10% | 69.34 / 74.50 | 62.63 / 77.33 | 60.40 / 77.89 | 61.55 / 76.92 |
+| 50% | 67.49 / 75.32 | 61.26 / 77.26 | 60.34 / 77.25 | 62.05 / 76.52 |
+| 100% | 65.95 / 75.77 | 61.18 / 76.97 | 62.17 / 76.39 | 62.43 / 76.31 |
+The results in Table˜24 clearly indicate that the hyperparameter exerts a more significant influence on the final OOD detection performance.
+Furthermore, a key relationship emerges: for optimal performance, a decrease in must be compensated by an increase in (and consequently, a higher ).
+This is explained by the formulation , where a smaller necessitates a larger value to ensure the term dominates the regularizer .
+| \ | 85 | 90 | 95 |
+| 0.1 | 70.07 / 74.12 | 87.45 / 66.06 | 91.81 / 56.09 |
+| 1 | 67.49 / 75.45 | 84.31 / 69.32 | 89.41 / 60.11 |
+| 10 | 58.97 / 78.98 | 69.47 / 79.68 | 78.77 / 76.26 |
+| 100 | 59.03 / 78.10 | 63.19 / 79.67 | 71.43 / 78.51 |
+In Table˜25, when is set to 0.1 with , does not dominate over term.
+As a result, it leads to scaling based on the regularization / balancing term instead of "predetermined" actual OOD detection signal "activation shift at peak activations".
+Such value of (e.g., ) invalidates the proposed hypothesis.
+From the results in Table˜26, it can be observed that activation shift with all activations considered () still proves to be a useful OOD detection signal.
+Furthermore, previous works (ASH, SCALE) observed that, particularly for ResNet-50 architecture, the performance drops after the percentile value of 85%.
+Similar observations can also be made in AdaSCALE from the results in Tables˜25 and 26.
+| \ | 85 | 90 | 95 |
+| 1% | 58.97 / 78.98 | 69.49 / 79.67 | 78.77 / 76.26 |
+| 5% | 60.36 / 77.74 | 68.30 / 78.23 | 78.19 / 75.58 |
+| 10% | 60.40 / 77.89 | 68.46 / 77.46 | 78.90 / 74.70 |
+| 50% | 60.34 / 77.25 | 69.88 / 76.39 | 80.46 / 73.17 |
+| 100% | 62.17 / 76.39 | 70.13 / 75.93 | 81.22 / 72.32 |
+Appendix F ISH regularization:
+Apart from enhancing the prior postprocessor ASH (Djurisic et al., 2023), SCALE (Xu et al., 2015) introduces a training regularization to emphasize samples with more distinct ID characteristics.
+We assess the performance (FPR@95 / AUROC ) of each method in ResNet-50 and ResNet-101 model following this regularization in Table˜27.
+The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by and in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively.
+Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in full-spectrum setting.
+Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%.
+These findings also generalize to ResNet-101 network.
+| Method | OOD Detection | FS-OOD Detection | ||
+| Near-OOD | Far-OOD | Near-OOD | Far-OOD | |
+| ResNet-50 | ||||
+| MSP | 74.07 / 62.16 | 51.13 / 84.64 | 87.52 / 40.36 | 74.41 / 61.52 |
+| MLS | 74.38 / 66.43 | 41.57 / 88.90 | 88.89 / 39.49 | 71.53 / 61.69 |
+| EBO | 74.68 / 66.46 | 41.85 / 88.83 | 89.05 / 39.18 | 71.77 / 61.11 |
+| ReAct | 71.98 / 70.81 | 28.76 / 93.49 | 87.78 / 43.88 | 61.87 / 71.64 |
+| ASH | 67.99 / 73.46 | 23.88 / 94.67 | 85.74 / 45.29 | 57.81 / 72.81 |
+| SCALE | 65.68 / 76.41 | 20.77 / 95.62 | 84.31 / 48.40 | 54.48 / 74.79 |
+| BFAct | 71.59 / 70.85 | 28.38 / 93.50 | 87.51 / 43.95 | 61.39 / 71.43 |
+| LTS | 66.32 / 75.03 | 22.07 / 95.28 | 85.08 / 46.16 | 57.11 / 73.06 |
+| OptFS | 67.71 / 73.03 | 24.65 / 94.18 | 85.38 / 45.91 | 57.09 / 72.95 |
+| AdaSCALE-A | 57.43 / 80.86 | 16.52 / 96.46 | 80.85 / 51.24 | 51.57 / 75.63 |
+| AdaSCALE-L | 56.83 / 80.81 | 17.62 / 96.22 | 80.97 / 50.59 | 53.43 / 74.62 |
+| ResNet-101 | ||||
+| MSP | 71.39 / 68.31 | 51.00 / 84.81 | 85.70 / 45.72 | 73.69 / 62.79 |
+| MLS | 72.32 / 71.94 | 41.04 / 88.99 | 87.39 / 44.88 | 69.94 / 63.23 |
+| EBO | 72.78 / 71.92 | 41.45 / 88.87 | 87.64 / 44.66 | 70.23 / 62.65 |
+| ReAct | 67.74 / 75.74 | 28.53 / 93.47 | 85.40 / 48.83 | 60.50 / 72.18 |
+| ASH | 66.03 / 77.79 | 25.21 / 94.43 | 83.95 / 50.57 | 56.91 / 73.37 |
+| SCALE | 64.30 / 78.98 | 23.09 / 94.95 | 83.14 / 51.05 | 55.88 / 73.27 |
+| BFAct | 67.53 / 75.86 | 28.32 / 93.40 | 85.11 / 48.96 | 60.12 / 71.83 |
+| LTS | 66.32 / 75.03 | 22.07 / 95.28 | 85.08 / 46.16 | 57.11 / 73.06 |
+| OptFS | 67.71 / 73.03 | 24.65 / 94.18 | 85.38 / 45.91 | 57.09 / 72.95 |
+| AdaSCALE-A | 54.66 / 83.52 | 16.81 / 96.32 | 78.52 / 55.19 | 49.92 / 76.20 |
+| AdaSCALE-L | 53.91 / 83.49 | 17.55 / 96.15 | 78.63 / 54.60 | 51.47 / 75.47 |
+Appendix G Hyperparameters
+| Dataset | Network | ||
+| CIFAR-10 | WideResNet-28-10 | 60 | 95 |
+| DenseNet-101 | 65 | 90 | |
+| CIFAR-100 | WideResNet-28-10 | 60 | 85 |
+| DenseNet-101 | 70 | 80 | |
+| ImageNet-1k | ResNet-50 | 80 | 85 |
+| ResNet-101 | 80 | 85 | |
+| RegNet-Y-16 | 60 | 90 | |
+| ResNeXt-50 | 80 | 85 | |
+| DenseNet-201 | 90 | 95 | |
+| EfficientNetV2-L | 60 | 99 | |
+| Vit-B-16 | 60 | 85 | |
+| Swin-B | 90 | 99 |
+| Dataset | Network | ||
+| CIFAR-10 | WideResNet-28-10 | 60 | 85 |
+| DenseNet-101 | 70 | 85 | |
+| CIFAR-100 | WideResNet-28-10 | 60 | 80 |
+| DenseNet-101 | 65 | 75 | |
+| ImageNet-1k | ResNet-50 | 80 | 85 |
+| ResNet-101 | 70 | 80 | |
+| RegNet-Y-16 | 60 | 85 | |
+| ResNeXt-50 | 70 | 80 | |
+| DenseNet-201 | 90 | 95 | |
+| EfficientNetV2-L | 60 | 99 | |
+| Vit-B-16 | 75 | 85 | |
+| Swin-B | 90 | 99 |
+Appendix H CIFAR-results
+H.1 WRN-28-10
 CIFAR-10 benchmark Method MNIST SVHN Textures Places365 Average MSP 17.02 / 94.61 21.71 / 92.96 60.50 / 88.06 42.27 / 90.04 35.38 / 91.42 MLS 13.01 / 96.76 30.35 / 93.06 76.12 / 86.65 52.56 / 90.46 43.01 / 91.73 EBO 12.93 / 96.93 30.35 / 93.12 76.15 / 86.68 52.57 / 90.56 43.00 / 91.82 ReAct 15.50 / 96.30 34.01 / 92.47 57.76 / 88.77 57.33 / 89.66 41.15 / 91.80 ASH 50.11 / 88.80 89.90 / 74.76 95.07 / 72.91 92.22 / 70.06 81.82 / 76.63 SCALE 13.24 / 96.70 32.21 / 92.88 75.76 / 86.77 55.81 / 90.09 44.26 / 91.61 BFAct 25.79 / 94.64 43.08 / 91.10 57.16 / 88.80 61.00 / 88.32 46.75 / 90.71 LTS 14.04 / 96.60 39.85 / 92.21 76.85 / 86.43 63.13 / 89.19 48.47 / 91.11 OptFS 25.68 / 94.83 51.58 / 89.86 62.14 / 88.07 80.19 / 84.05 54.90 / 89.20 AdaSCALE-A 14.93 / 96.02 17.84 / 95.14 64.96 / 88.31 34.57 / 92.31 33.08 / 92.95 AdaSCALE-L 15.58 / 95.98 18.41 / 95.10 62.87 / 88.67 37.59 / 91.97 33.61 / 92.93 CIFAR-100 benchmark MNIST SVHN Textures Places365 Average MSP 49.79 / 78.72 56.76 / 80.70 64.49 / 76.86 56.66 / 79.96 56.92 / 79.06 MLS 46.57 / 81.43 53.08 / 83.37 64.59 / 77.65 59.70 / 79.82 55.99 / 80.57 EBO 46.41 / 81.99 52.92 / 83.77 64.58 / 77.61 59.76 / 79.60 55.92 / 80.74 ReAct 49.92 / 81.07 40.66 / 86.49 52.42 / 80.81 60.35 / 79.72 50.84 / 82.03 ASH 44.06 / 85.55 41.48 / 87.50 61.78 / 81.65 80.45 / 71.83 56.94 / 81.63 SCALE 40.65 / 84.68 48.56 / 85.56 58.45 / 80.81 60.51 / 79.90 52.04 / 82.74 BFAct 61.59 / 77.47 34.74 / 88.50 47.30 / 83.38 64.49 / 78.47 52.03 / 81.96 LTS 36.27 / 87.38 45.41 / 87.23 53.90 / 83.18 62.62 / 79.64 49.55 / 84.36 OptFS 57.61 / 79.47 37.04 / 86.43 53.02 / 80.43 70.44 / 76.77 54.53 / 80.78 AdaSCALE-A 45.18 / 81.69 36.79 / 89.20 55.93 / 81.93 56.48 / 81.55 48.59 / 83.59 AdaSCALE-L 42.13 / 83.58 32.44 / 91.02 50.87 / 84.14 57.83 / 81.51 45.82 / 85.06
 Method CIFAR-10 benchmark CIFAR-100 benchmark Average CIFAR-100 TIN CIFAR-10 TIN MSP 54.13 / 88.28 42.94 / 89.93 56.83 / 80.42 48.82 / 83.39 50.68 / 85.51 MLS 67.10 / 87.72 55.91 / 89.90 58.99 / 80.98 49.27 / 84.01 57.82 / 85.65 EBO 67.04 / 87.77 55.88 / 89.97 58.97 / 80.93 49.39 / 83.95 57.82 / 85.66 ReAct 65.96 / 87.76 51.44 / 90.33 69.17 / 79.10 51.56 / 83.77 59.53 / 85.24 ASH 91.33 / 70.72 90.77 / 73.28 85.25 / 69.96 78.31 / 74.55 86.42 / 72.13 SCALE 69.52 / 87.35 59.48 / 89.53 61.30 / 80.35 51.25 / 83.65 60.39 / 85.22 BFAct 66.31 / 86.94 57.12 / 89.69 78.90 / 74.98 59.04 / 82.25 65.34 / 83.47 LTS 74.28 / 86.38 66.01 / 88.56 64.17 / 79.57 54.24 / 83.09 64.68 / 84.40 OptFS 76.36 / 84.05 66.73 / 86.56 85.40 / 75.55 64.11 / 80.99 73.15 / 79.83 AdaSCALE-A 50.60 / 89.40 42.80 / 91.13 62.21 / 79.99 47.11 / 84.98 50.68 / 86.38 AdaSCALE-L 53.98 / 89.01 45.95 / 90.77 65.41 / 79.27 48.74 / 84.75 53.52 / 85.95
-D.2 DenseNet-101
+H.2 DenseNet-101
 CIFAR-10 benchmark Method MNIST SVHN Textures Places365 Average MSP 17.91 / 94.22 32.04 / 90.38 46.80 / 87.53 37.59 / 89.24 33.59 / 90.34 MLS 10.02 / 97.58 31.25 / 92.59 64.43 / 85.58 39.19 / 90.74 36.22 / 91.62 EBO 9.74 / 97.76 31.23 / 92.69 64.46 / 85.48 39.17 / 90.81 36.15 / 91.68 ReAct 12.60 / 97.24 34.79 / 92.02 50.41 / 88.21 36.12 / 91.35 33.48 / 92.20 ASH 9.40 / 98.12 39.42 / 91.25 70.95 / 85.39 57.90 / 85.48 44.42 / 90.06 SCALE 9.04 / 97.88 26.99 / 93.54 61.52 / 86.77 39.57 / 90.76 34.28 / 92.24 BFAct 23.59 / 94.96 42.49 / 89.00 53.74 / 87.38 37.53 / 91.09 39.34 / 90.61 LTS 8.92 / 97.97 27.16 / 93.59 59.07 / 87.09 39.47 / 90.81 33.65 / 92.37 OptFS 9.74 / 97.88 41.20 / 90.71 51.35 / 88.48 59.47 / 86.03 40.44 / 90.77 AdaSCALE-A 12.42 / 96.85 25.04 / 94.05 58.28 / 87.35 36.77 / 91.20 33.13 / 92.36 AdaSCALE-L 10.92 / 97.44 26.43 / 93.87 58.59 / 87.19 37.03 / 91.25 33.24 / 92.44 CIFAR-100 benchmark MSP 65.65 / 72.43 63.81 / 76.52 75.34 / 72.19 61.36 / 77.16 66.54 / 74.57 MLS 58.69 / 78.55 57.12 / 79.43 79.05 / 72.68 62.72 / 78.38 64.39 / 77.26 EBO 58.58 / 78.98 56.76 / 79.19 79.09 / 72.44 62.86 / 78.08 64.32 / 77.17 ReAct 62.71 / 76.37 48.48 / 81.64 64.65 / 78.62 59.00 / 78.89 58.71 / 78.88 ASH 40.69 / 88.57 48.03 / 86.24 65.24 / 83.29 73.29 / 71.88 56.81 / 82.49 SCALE 56.92 / 79.53 53.81 / 80.96 76.10 / 74.47 62.49 / 78.51 62.33 / 78.37 BFAct 73.83 / 67.19 60.01 / 75.85 69.29 / 76.41 68.15 / 73.73 67.82 / 73.29 LTS 55.33 / 80.58 51.33 / 82.04 73.06 / 75.89 62.58 / 78.36 60.58 / 79.22 OptFS 64.24 / 75.24 59.81 / 76.46 66.15 / 77.50 73.47 / 69.76 65.92 / 74.74 AdaSCALE-A 62.51 / 74.96 46.29 / 84.31 71.40 / 76.59 61.70 / 78.86 60.47 / 78.68 AdaSCALE-L 61.33 / 75.73 43.97 / 85.30 69.31 / 77.71 61.97 / 78.69 59.15 / 79.36
 Method CIFAR-10 benchmark CIFAR-100 benchmark Average CIFAR-100 TIN CIFAR-10 TIN MSP 40.13 / 88.45 35.50 / 89.61 59.94 / 77.53 56.96 / 79.57 48.13 / 83.79 MLS 45.14 / 88.85 38.01 / 90.85 63.61 / 78.26 57.09 / 81.75 50.96 / 84.93 EBO 45.19 / 88.85 38.05 / 90.90 63.90 / 77.94 57.53 / 81.58 51.17 / 84.82 ReAct 44.34 / 89.19 37.10 / 91.08 70.77 / 75.06 61.30 / 80.40 53.38 / 83.93 ASH 67.78 / 82.68 62.54 / 85.18 81.65 / 65.84 78.66 / 70.01 72.66 / 75.93 SCALE 45.25 / 88.92 37.76 / 91.01 64.20 / 78.13 56.96 / 81.88 51.04 / 84.99 BFAct 52.06 / 87.70 44.09 / 89.89 79.31 / 67.39 71.79 / 74.18 61.81 / 79.79 LTS 44.89 / 89.03 37.67 / 91.10 64.66 / 77.87 56.83 / 81.87 51.01 / 84.97 OptFS 60.63 / 85.29 55.55 / 86.96 82.97 / 64.99 74.73 / 70.55 68.47 / 76.95 AdaSCALE-A 43.29 / 89.37 35.57 / 91.32 65.51 / 78.00 54.49 / 82.44 49.72 / 85.28 AdaSCALE-L 43.19 / 89.40 35.70 / 91.37 66.09 / 77.79 54.57 / 82.47 49.89 / 85.26
-Appendix E ImageNet-1k results
-E.1 near-OOD detection
+Appendix I ImageNet-1k results
+I.1 near-OOD detection
 Method SSB-Hard NINCO ImageNet-O Average MSP 74.49 / 72.09 56.88 / 79.95 91.32 / 28.60 74.23 / 60.21 MLS 76.20 / 72.51 59.44 / 80.41 88.97 / 40.73 74.87 / 64.55 EBO 76.54 / 72.08 60.58 / 79.70 88.84 / 41.78 75.32 / 64.52 REACT 77.55 / 73.03 55.82 / 81.73 84.45 / 51.67 72.61 / 68.81 ASH 73.66 / 72.89 53.05 / 83.45 81.70 / 57.67 69.47 / 71.33 SCALE 67.72 / 77.35 51.80 / 85.37 83.77 / 59.89 67.76 / 74.20 BFAct 77.20 / 73.15 55.27 / 81.88 84.57 / 51.62 72.35 / 68.88 LTS 68.46 / 77.10 51.24 / 85.33 84.33 / 57.69 68.01 / 73.37 OptFS 78.32 / 71.01 52.09 / 82.51 78.56 / 59.40 69.66 / 70.97 AdaSCALE-A 57.96 / 81.68 44.92 / 87.15 74.06 / 68.12 58.98 / 78.98 AdaSCALE-L 58.68 / 81.42 45.01 / 87.11 75.83 / 67.33 59.84 / 78.62
@@
 | AdaSCALE-L | 80.12 / 70.06 | 63.68 / 81.35 | 74.87 / 72.34 | 72.89 / 74.58 |
-E.2 far-OOD detection
+I.2 far-OOD detection
 | Method | iNaturalist | Textures | OpenImage-O | Places | Average |
@@
 | AdaSCALE-L | 30.95 / 91.69 | 60.17 / 86.30 | 37.52 / 90.08 | 56.32 / 83.82 | 46.24 / 87.97 |
-E.3 Full-Spectrum near-OOD detection
+I.3 Full-Spectrum near-OOD detection
 | Method | SSB-Hard | NINCO | ImageNet-O | Average |
@@
 | AdaSCALE-L | 90.18 / 46.63 | 80.77 / 57.85 | 87.27 / 47.90 | 86.07 / 50.79 |
-E.4 Full-Spectrum far-OOD detection
+I.4 Full-Spectrum far-OOD detection
 | Method | iNaturalist | Textures | OpenImage-O | Places | Average |